### v1

# Positioning-Aided Channel Estimation for Multi-LEO Satellite Downlink Communications

Abstract
We investigate a multi-low Earth orbit (LEO) satellite system that simultaneously provides positioning and communication services to terrestrial user terminals. To address the challenges of channel estimation in LEO satellite systems, we propose a novel two-timescale positioning-aided channel estimation framework, exploiting the distinct variation rates of position-related parameters and channel gains inherent in LEO satellite channels. Using the misspecified Cramér-Rao bound (MCRB) theory, we systematically analyze positioning performance under practical imperfections, such as inter-satellite clock bias and carrier frequency offset. Furthermore, we theoretically demonstrate how position information derived from downlink positioning can enhance uplink channel estimation accuracy, even in the presence of positioning errors, through an MCRB-based analysis. To overcome the constraints of limited link budgets and communication rates associated with single-satellite-based communication, we develop a distributed beamforming strategy for downlink communication. This strategy allows LEO satellites to independently optimize their beamformers using local channel state information, eliminating the need for centralized processing while preserving the advantages of multi-satellite cooperative communication. Theoretical analyses and numerical results confirm the effectiveness of the proposed framework in achieving high-precision downlink positioning under practical imperfections, facilitating uplink channel estimation, and enabling efficient downlink communication.
Index Terms:
LEO satellite positioning and communication, channel estimation, MCRB, distributed beamforming.
I Introduction
The evolution of mobile communications has been driven by a remarkable vision: enabling seamless, high-speed information access from any point on Earth. While generational advances in cellular networks, from 1G to 5G and beyond, have largely achieved this goal in urban and suburban areas where infrastructure deployment is economically viable, significant challenges remain in rural and remote regions. In these underdeveloped areas, reliable communication and internet access often remain either unavailable or prohibitively expensive, contributing to a growing digital divide[1]. The non-terrestrial networks (NTN)-based communications have emerged as a promising solution to bridge this gap, with airborne and spaceborne platforms such as high-altitude platform stations and satellites, serving as access points to enable truly global connectivity[2]. Among these candidates, low Earth orbit (LEO) satellites have attracted particular attention due to their distinct advantages over medium Earth orbit (MEO) and geostationary Earth orbit (GEO) satellites, including reduced path loss and propagation delay, greater constellation flexibility, and lower launch costs[3, 4, 5].
The advancement of LEO satellite-based global connectivity presents promising opportunities alongside significant challenges, particularly in channel estimation, a crucial aspect for meeting increasingly demanding high-speed communication requirements[6, 7]. While receivers can effectively mitigate channel aging by compensating for Doppler shift caused by rapid LEO satellite movement[3, 6, 8], the substantial propagation distance compared to terrestrial networks results in severe signal power attenuation. This attenuation particularly impacts pilot-based channel estimation methods, which struggle with low receive power, making effective channel state information (CSI) acquisition challenging, especially for power-constrained user terminals such as mobile devices in the uplink. On the other hand, despite the reduced orbital altitude compared to MEO and GEO satellites, LEO communication links still span distances vastly exceeding typical terrestrial communication counterparts. This extended signal travel distance, combined with limited onboard power resources, significantly constrains the link budget of single-LEO satellite-based communication systems, creating a substantial barrier to achieving the ambitious throughput goals of wideband NTN communications[2, 9, 7].
Although satellite channels pose significant challenges for CSI acquisition, they exhibit distinct characteristics compared to terrestrial channels, particularly stronger line-of-sight (LOS) dominance and geometric properties [9]. These characteristics are inherently related to position information [5], suggesting that strategic utilization of UT position information could potentially address channel estimation challenges in LEO satellite communications, especially by enabling partial channel inference through the reconstruction of components such as steering vectors. However, this approach raises critical research questions, such as how to achieve accurate and efficient position acquisition within LEO satellite communication networks and how to effectively leverage this position information to enhance LEO satellite channel estimation. On the other hand, to address the limited link budget of single-satellite systems, multi-satellite cooperative transmission has emerged as a promising solution [10, 11, 12, 13, 14, 15]. However, achieving coherent transmission across satellites within non-orthogonal frequency or time resources presents significant challenges, such as phase synchronization among different satellites and the need for centralized processing based on global CSI[10]. These centralized operations introduce substantial processing overhead and latency, further complicating the practical implementation of cooperative transmission schemes.
The role of position information in terrestrial networks has been widely recognized for enhancing communication performance[16, 17, 18, 19, 20]. For instance, [16] introduces location awareness in 5G networks, demonstrating its potential for synchronizing coordinated communication schemes. Similarly, [17] showcases how position information facilitates beam alignment in millimeter-wave communication. These approaches typically rely on position information from either global navigation satellite system (GNSS)[16, 17] or radio sensing [18, 19, 20]. However, in LEO satellite scenarios, obtaining position information of terrestrial UTs through radar-like sensing at LEO satellites becomes impractical due to excessive round-trip signal attenuation. While GNSS appears to be a viable alternative, as demonstrated in [21], where UTs’ positions are obtained via GNSS and fed back to LEO satellites for downlink communication, the critical impact of positioning errors and the reliance on overidealistic synchronization assumptions remains under-explored and warrant further investigation in practical implementations.
Concurrent with advances in LEO satellite communication, there is growing interest in exploring LEO satellites as potential supplements or alternatives to GNSS [22, 23]. This interest arises from the inherent advantages of LEO satellites over GNSS, including stronger signal reception, extensive constellations, and broad frequency diversity, characteristics that have also partially driven the success of LEO satellite communications [23, 24, 25]. Recent research has begun to investigate the integration of LEO satellite communication and positioning functionalities. For instance, [26] proposes a massive multi-input-multi-output (MIMO) LEO satellite system that supports simultaneous communication and positioning in the downlink, designing hybrid beamforming to meet their distinct requirements. However, this work primarily focuses on the tradeoff between positioning and communication performance, overlooking their mutual benefits such as the potential of positioning in facilitating channel estimation, which is a critical challenge in LEO satellite systems. Moreover, it does not account for the inevitable clock bias and carrier frequency offset (CFO) among LEO satellites, factors that may significantly affect positioning performance.
Three key questions emerge under this context: 1) how can the positioning performance of LEO satellites be effectively characterized under practical imperfections such as clock bias and CFO; 2) how can the position information of UTs obtained through LEO satellite positioning be leveraged to assist the challenging channel estimation task; and 3) how can the benefits of the improved link budget offered by multi-satellite cooperative communication be harnessed while avoiding the complexities of synchronization and centralized processing?
In this paper, we investigate a scenario where a cluster of LEO satellites, each equipped with a uniform planar array (UPA), simultaneously provides downlink positioning and communication services to multiple single-antenna terrestrial UTs. We aim to: 1) characterize the positioning performance under practical impairments such as clock bias/CFO incurred by imperfect time-/frequency-domain synchronization; 2) unveil how the UTs’ position information obtained by downlink positioning can be exploited to enhance uplink channel estimation; and 3) explore distributed beamforming relying solely on local CSI acquired at each satellite, enabling efficient downlink communication while avoiding the overhead and complexity typically associated with centralized operations. The main contributions of this paper are as follows.
  • Building on the distinct variation rates of position-related parameters and random channel gains in LEO satellite channels, as detailed in Section II, we propose a two-timescale frame structure. Downlink positioning operates on the longer position-coherent timescale, while uplink channel estimation and downlink communication occur on the shorter channel gain-coherent timescale. Through this framework, each process is scheduled according to its variation rate, maximizing resource utilization while preserving the potential to leverage positioning for efficient channel estimation and communication.
  • To characterize the LEO satellite-based positioning performance under imperfect inter-satellite clock bias and CFO, we employ misspecified Cramér-Rao bound (MCRB) theory rather than traditional Cramér-Rao bound (CRB) analysis. The MCRB framework is specifically designed to derive performance limits for estimation under model mismatch[27]. Through this theoretical approach, we rigorously analyze how imperfect inter-satellite clock bias and CFO affect time-of-arrival (TOA) and Doppler shift measurements, and systematically derive their impact on positioning accuracy.
  • We demonstrate the advantage of leveraging the position information of UTs, obtained during the positioning stage, to enhance uplink channel estimation. Specifically, we account for the inevitable positioning errors, which introduce model mismatches in the positioning-aided channel estimation (PACE) process. By invoking the MCRB theory, we provide a theoretical characterization of the uplink channel estimation performance facilitated by the UTs’ position information.
  • We propose a distributed beamforming strategy for downlink communication, where each UT is served by the LEO satellites withn the cluster simultaneously. In this approach, each LEO satellite independently optimizes its beamformer using only its local CSI with respect to UTs, thereby eliminating the need for a centralized processor and complex inter-satellite coordination mechanisms.
This paper is structured as follows. We begin by presenting the system model in Section II. Section III examines both downlink positioning and uplink channel estimation performance through MCRB theory, incorporating the PACE. A distributed beamforming approach for downlink communications is detailed in Section IV. Numerical results are presented in Section V, with conclusions drawn in Section VII.
Notations: Scalars are written in regular lowercase, while vectors and matrices are denoted by bold lowercase and bold uppercase letters, respectively. For a vector $\mathbf{a}$, its 2-norm is expressed as $\left\|\mathbf{a}\right\|$. We use $\mathsf{T}$ and $\mathsf{H}$ as superscripts to indicate transpose and Hermitian transpose operations. For a matrix $\mathbf{A}$, $\text{Tr}(\mathbf{A})$ represents its trace. Complex numbers are handled by $\Re\left\{a\right\}$ and $\Im\left\{a\right\}$ for their real and imaginary components. A block-diagonal matrix constructed from matrices $\mathbf{A}_{1},\ldots,\mathbf{A}_{N}$ is written as $\text{blkdiag}\{\mathbf{A}_{1},\ldots,\mathbf{A}_{N}\}$. The Kronecker product is symbolized by $\otimes$. For statistical distributions, $\mathcal{CN}(\bm{\mu},\mathbf{C})$ represents a circularly symmetric complex Gaussian distribution characterized by mean $\bm{\mu}$ and covariance matrix $\mathbf{C}$.
II System Model
Figure 1: A cluster of LEO satellites traverses orbit while simultaneously delivering positioning and communication services to UTs. Each UT is served by all LEO satellites within the cluster for both functionalities. Positioning operations in the downlink occur over a relatively large timescale, whereas uplink channel estimation and downlink communication are carried out on a smaller timescale. The specific definitions of these timescales will be clarified later.
As illustrated in Fig. 1, we consider a scenario where a cluster of $S$ LEO satellites serves $U$ UTs, simultaneously providing downlink positioning and communication services. Each LEO satellite is equipped with a UPA comprising $N=N_{\text{h}}\times N_{\text{v}}$ half-wavelength-spaced antennas, where $N_{\text{h}}$ and $N_{\text{v}}$ denote the numbers of antennas in the horizontal and vertical dimensions, respectively. Each UT is equipped with a single antenna. While accurate CSI is essential for achieving efficient downlink communication, channel estimation is challenging in space-borne communication systems due to the limited link budget and short coherence time. These challenges originate from the long travel distance between the satellites and the UTs as well as the high-speed satellite movement[7]. In the following, we first introduce the adopted channel model. Subsequently, we elaborate on a two-timescale property of this model, which can be leveraged to facilitate efficient channel acquisition at LEO satellites, thereby enhancing downlink communications.
II-A Channel Model
We consider a satellite communication system operating at relatively high-frequency bands such as the Ku and Ka bands, where the channel is dominated by LOS propagation[9, 13]. [^1] While the multipath effect can still play a significant role in dense urban environments, LEO satellite communications are primarily designed to serve suburban, rural, and remote areas, such as oceans and deserts, where terrestrial infrastructure is sparse or nonexistent, and scatterers are greatly reduced. Consequently, LOS propagation becomes even more dominant. Assume the system employs the orthogonal frequency division multiplexing (OFDM) scheme with $K$ subcarriers. Let $\Delta f=B/K$ and $T=1/B$ denote the subcarrier spacing and symbol duration, respectively, where $B$ is the signal bandwidth. The channel from the $s$-th LEO satellite to the $u$-th UT during the $\ell$-th symbol over the $k$-th subcarrier is expressed as [^2] Note that, though not further indexed for symbolic simplicity, this expression is viewed within one channel-gain coherent interval (as elaborated later) and $\upsilon_{s,u}$, $\tau_{s,u}$, and $\bm{\theta}_{s,u}$ are viewed constant for many such intervals.
$\mathbf{h}_{s,u}\left[\ell,k\right]=\alpha_{s,u}G\left(\theta_{s,u}^{\mathrm{el}}\right)e^{\jmath 2\pi\left(\ell T\upsilon_{s,u}-k\Delta f\tau_{s,u}\right)}\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (1)
where $\alpha_{s,u}$ is the complex channel gain, $G(\cdot)$ is the antenna radiation gain determined solely by the elevation angle[28], $\tau_{s,u}$ and $\upsilon_{s,u}$ represent the TOA and Doppler shift, respectively, and $\mathbf{a}(\bm{\theta}_{s,u})\in\mathbb{C}^{N}$ is the steering vector at the LEO satellite, with $\bm{\theta}_{s,u}=[\theta_{s,u}^{\mathrm{az}},\theta_{s,u}^{\mathrm{el}}]^{\mathsf{T}}$ being the angle-of-departure (AOD) composed of both azimuth and elevation angles. Without loss of generality, we assume the UPA at each satellite is deployed on the XY-plane of its local coordinate system, as illustrated in Fig. 2. Hence, the steering vector can be expressed as
$\mathbf{a}\left(\bm{\theta}_{s,u}\right)=e^{-\jmath 2\pi\phi_{s,u}^{\text{h}}\mathbf{n}\left(N_{\text{h}}\right)}\otimes e^{-\jmath 2\pi\phi_{s,u}^{\text{v}}\mathbf{n}\left(N_{\text{v}}\right)},$ (2)
where $\mathbf{n}(N)=[0,\ldots,N-1]^{\mathsf{T}}$, $\phi_{s,u}^{\text{h}}=d\cos\theta_{s,u}^{\mathrm{az}}\cos\theta_{s,u}^{\mathrm{el}}/\lambda$, and $\phi_{s,u}^{\text{v}}=d\sin\theta_{s,u}^{\mathrm{az}}\cos\theta_{s,u}^{\mathrm{el}}/\lambda$. Here, $d$ denotes the antenna spacing along each dimension, and $\lambda$ is the wavelength at the central carrier frequency $f_{c}$.
The complex channel gain can be modeled as
$\alpha_{s,u}=e^{\jmath\psi_{s,u}}\beta_{s,u},$ (3)
where $\psi_{s,u}$ represents the random phase difference between the transmitter and receiver [29], and $\beta_{s,u}$ denotes the path loss incorporating both the large-scale path losses and small-scale fading, as described in [30]. Specifically, the path loss $\beta_{s,u}$, in the dB domain, is characterized as
[…]
where $\beta^{\text{FS}}_{s,u}$ denotes the free-space path loss, $\beta^{\text{SF}}_{s,u}$ accounts for shadow fading modeled as a Gaussian random variable, $\beta^{\text{CL}}_{s,u}$ represents the clutter loss, $\beta^{\text{AB}}_{s,u}$ captures atmospheric absorption effects, and $\beta^{\text{SC}}_{s,u}$ accounts for attenuation caused by ionospheric or tropospheric scintillation.
Figure 2: Geometric configuration of the local coordinate systems for LEO satellites, depicting the UPA deployment in the XY-plane.
II-B Timescale Analysis and Frame Structure
II-B1 Timescale Analysis
From (1), we observe that the LEO satellite channel is governed by two categories of parameters:
  • Large-Timescale Parameters: These parameters are associated with the UT’s position, including the Doppler shift, TOA, and AOD. Due to the large geographical scale of satellite communications, they typically change slowly (on timescales of seconds) or can be accurately predicted through tracking mechanisms to reduce estimation overhead[25]. Although LEO satellites move rapidly, their orbits are typically predictable and well-determined. Therefore, while these parameters are related to the fast-varying position of LEO satellites, they can still be considered slow-varying due to their predictability.
  • Small-Timescale Parameters: These parameters consist of random variations in the amplitude and phase of the channel gain. In contrast to the slow-varying parameters determined by geometry, these parameters are inherently unpredictable and fluctuate rapidly. For instance, tropospheric scintillation can cause phase and amplitude variations in Ka-band LEO satellite channels that fluctuate on sub-second timescales, with the exact rate depending on atmospheric and orbital conditions[31].
Notably, the Doppler shift caused by the high-speed movement of the satellite induces rapid temporal phase rotation in the channel. However, this effect can be largely mitigated through carrier synchronization at the receiver, albeit with minor residual errors [3]. As a result, its impact on LEO satellite communications is limited. Nonetheless, we retain it in the channel model because it can serve as a valuable measurement for positioning, as will be discussed later. For positioning purposes, the Doppler shift should be explicitly estimated before being compensated for communications.
II-B2 Two-Timescale Frame Structure
Figure 3: Illustration of the two-timescale frame structure, where each large-timescale frame encompasses multiple small-timescale subframes. The large-timescale frame is governed by the position-coherent time, while the small-timescale subframe is defined by the channel gain-coherent time.
Inspired by the timescale analysis, we propose a two-timescale frame structure that exploits the inherent two-timescale property of LEO channels to enable efficient channel estimation/reconstruction, thus facilitating downlink LEO transmission. As shown in Fig. 3, the duration of a large-timescale frame is determined by the position-coherent time, which indicates the interval after which the UTs’ position information becomes outdated. Each large-timescale frame begins with an initial subframe for downlink positioning (where position estimates obtained by UTs are fed back to LEO satellites), followed by multiple small-timescale subframes during which the channel gain remains coherent. Each small-timescale subframe is further divided into slots for uplink channel estimation and downlink communication. Note that time division duplex (TDD) is adopted to differentiate between uplink and downlink transmissions. The proposed approach leverages uplink-downlink channel reciprocity, which applies uplink channel estimation to optimize downlink communication design.
To further substantiate the fundamental rationales behind the proposed structure for LEO downlink communications, we highlight its potential advantages by comparing it with positioning-free uplink channel estimation protocols, where the channel is directly estimated at the LEO satellites using uplink pilots without requiring position information from UTs[6, 8]. This approach involves estimating an unknown channel vector/matrix, which can incur significant pilot overhead, especially when large antenna arrays are employed. With the proposed position information-aided framework, vector-type channel estimation can be transformed into scalar-type estimation by exploiting the channel structure. This transformation offers the potential for higher channel estimation accuracy with the same pilot overhead. Alternatively, to achieve equivalent channel estimation accuracy, reduced pilot overhead is required, thus preserving more resources for data transmission.
In the following, we will first conduct a performance analysis based on the two-timescale PACE and then propose a distributed beamforming method across multiple LEO satellites for downlink communication.
III Two-timescale Positioning-Aided Channel Estimation
To characterize the performance limits of the proposed two-timescale PACE, we conduct a two-phase theoretical performance analysis. In the first phase, we analyze the downlink multi-LEO satellite-based positioning performance, while in the second phase, we assess the uplink channel estimation performance.
Before analyzing positioning performance, we acknowledge the presence of model mismatches due to hardware imperfections. First, low-cost oscillators and inherent hardware defects can introduce phase and frequency errors [32]. Second, clock biases are inevitable between satellites and UTs as well as between satellites themselves. Although a sub-nanosecond clock synchronization can be achieved by deploying GNSS receivers on LEO satellites [33], actual satellite signals remain subject to ionospheric and tropospheric delays during propagation [34]. These delays, influenced by factors such as temperature, pressure, humidity, and the satellite’s elevation angle relative to the UT, result in unique delays for each satellite-UT pair, effectively introducing an additional clock bias in the signal model. Despite these non-idealities, many localization algorithms simplify the model by ignoring such factors to reduce computational complexity. For instance, in the scenario considered here, the large beam footprint of LEO satellite communication systems relaxes UT positioning accuracy requirements, while rapid channel variations prioritize model simplicity for timely estimations. Accordingly, we adopt a simplified model to estimate unknowns derived from a more complex realistic model, which inherently leads to model mismatches. Similarly, in the second phase, namely position information-aided channel estimation, the presence of estimation errors in UT positions also introduces model mismatches. Consequently, the performance bound under these model mismatches can be effectively analyzed using the tool of MCRB [27].
III-A MCRB Basics
This subsection briefly recaps the fundamentals of MCRB. For a parameter $\bar{\bm{\theta}}\in\mathbb{R}^{M\times 1}$, the lower bound (LB) matrix of the mean squared error of a mismatched estimator is provided by[27]
$\text{LBM}\big{(}\tilde{\bm{\theta}}\big{)}=\underbrace{\mathbf{A}^{-1}_{\tilde{\bm{\theta}}}\mathbf{B}_{\tilde{\bm{\theta}}}\mathbf{A}^{-1}_{\tilde{\bm{\theta}}}}_{\text{MCRB}\left(\tilde{\bm{\theta}}\right)}+\underbrace{\left(\bar{\bm{\theta}}-\tilde{\bm{\theta}}\right)\left(\bar{\bm{\theta}}-\tilde{\bm{\theta}}\right)^{\top}}_{\text{Bias}\left(\tilde{\bm{\theta}}\right)}.$ (5)
Here, $\tilde{\bm{\theta}}$ denotes the pseudo-true parameter obtained by
$\tilde{\bm{\theta}}=\arg\underset{\bm{\theta}}{\min\;}\mathcal{D}\left\{\xi_{\text{T}}\left(\mathbf{y};\bar{\bm{\theta}}\right)\left|\right|\xi_{\text{M}}\left(\mathbf{y};\bm{\theta}\right)\right\},$ (6)
where $\mathcal{D}\{\xi_{\text{T}}\left(\mathbf{y};\bm{\theta}\right)\left|\right|\xi_{\text{M}}(\mathbf{y};\tilde{\bm{\theta}})\}$ denotes the Kullback-Leibler (KL) divergence between the true likelihood function $\xi_{\text{T}}\left(\mathbf{y};\bm{\theta}\right)$ and the mismatched likelihood function $\xi_{\text{M}}(\mathbf{y};\tilde{\bm{\theta}})$. In addition, $\mathbf{A}_{\tilde{\bm{\theta}}}$ and $\mathbf{B}_{\tilde{\bm{\theta}}}$ represent two generalized Fisher information matrixs, whose elements in the $i$-th row and the $j$-th column are determined by[27]
$\left[\mathbf{A}_{\tilde{\bm{\theta}}}\right]_{i,j}=$ $\mathbb{E}_{f_{T}}\left\{\frac{\partial^{2}\ln\xi_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}\partial\left[\bm{\theta}\right]_{j}}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}}\right\}$
$=$ $2\Re\left\{\left(\frac{\partial^{2}\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}\partial\left[\bm{\theta}\right]_{j}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\bm{\epsilon}\left(\bm{\theta}\right)\right.$ (7)
$\;\;\;\;\;\;\;\;\;-\left.\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right)\right\}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}},$
and
$\left[\mathbf{B}_{\tilde{\bm{\theta}}}\right]_{i,j}=$ $\mathbb{E}_{f_{T}}\left\{\frac{\partial\ln f_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}}\frac{\partial\ln f_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}}\right\}$
$=$ $4\Re\left\{\bm{\epsilon}\left(\bm{\theta}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right\}\Re\left\{\bm{\epsilon}\left(\bm{\theta}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right\}$
$+$ $2\Re\left\{\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right)\right\}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}},$ (8)
respectively. Here, $\mathbf{C}_{\text{M}}$ represents the covariance matrix of $\xi_{\text{M}}(\mathbf{y};\bm{\theta})$ and $\bm{\epsilon}(\tilde{\bm{\theta}})=\bm{\kappa}(\bar{\bm{\theta}})-\bm{\mu}(\tilde{\bm{\theta}})$ with $\bm{\kappa}(\bar{\bm{\theta}})$ and $\bm{\mu}(\tilde{\bm{\theta}})$ being the noise-free observations under the true and mismatched models, respectively[27, 35].
III-B Large-Timescale Downlink Positioning
III-B1 Signal Model
Let $\mathbf{t}_{s}[\ell,k]\in\mathbb{C}^{U}$ be the pilot (with unit-modulus elements) sent by the $s$-th LEO satellite over the $\ell$-th symbol and the $k$-th subcarrier. We assume the satellites are multiplexed in the positioning phase through time-division multiple access (TDMA)[25]. The received baseband signal at the $u$-th user, from the $s$-th satellite can be expressed as
$y_{s,u}\left[\ell,k\right]=$ $\mathbf{h}_{s,u}^{\mathsf{T}}\left[\ell,k\right]\mathbf{F}_{s}\mathbf{t}_{s}\left[\ell,k\right]+z_{u}\left[\ell,k\right]$
$=$ $\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}^{\mathsf{T}}\left(\bm{\theta}_{s,u}\right)\mathbf{f}_{s}\left[\ell,k\right]+z_{u}\left[\ell,k\right],$ (9)
[…]
Let ${\bm{\eta}}_{s,u}=[\acute{\bm{\eta}}_{s,u}^{\mathsf{T}},\grave{\bm{\eta}}_{s,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{6}$ be the unknown channel-domain parameters in signals received from the $s$-th satellite, where $\acute{\bm{\eta}}_{s,u}=[\upsilon_{s,u},\tau_{s,u}]^{\mathsf{T}}\in\mathbb{R}^{2}$ contains the channel-domain parameters we used to localize the user and $\grave{\bm{\eta}}_{s,u}=[\bm{\theta}_{s,u}^{\mathsf{T}},\Re(\alpha_{s,u}),\Im(\alpha_{s,u})]^{\mathsf{T}}\in\mathbb{R}^{4}$ contains the other nuisance parameters that we do not explore. Note here the AOD $\bm{\theta}_{s,u}$ is unused for UT positioning due to the limited angular resolution of satellites. We further concatenate all the utilized channel-domain parameters collected from $S$ satellites as $\acute{\bm{\eta}}_{u}=[\acute{\bm{\eta}}_{1,u}^{\mathsf{T}},\ldots,\acute{\bm{\eta}}_{S,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{2S}$. Therefore, in the large-timescale downlink positioning, we first estimate $\acute{\bm{\eta}}_{u}$ based on $y_{s,u}[\ell,k]$, $s=1,\dots,S$, $\ell=1,\dots,L_{\text{p}}$, $k=1,\dots,K$, where $L_{\text{p}}$ denotes the number of pilot symbols for downlink positioning. Afterwards, we estimate the $u$-th user’s position, clock biases, and CFOs.
III-B2 The FIM of ${\bm{\eta}}_{s,u}$
Let ${\mu}_{s,u}[\ell,k]$ denote the noise-free version of $y_{s,u}[\ell,k]$ and $\mathbf{G}_{s,u}[\ell,k]=[\mathbf{g}_{v}[\ell,k],\mathbf{g}_{\tau}[\ell,k],\mathbf{g}_{\theta_{\mathrm{az}}}[\ell,k],\mathbf{g}_{\theta_{\mathrm{el}}}[\ell,k],\mathbf{g}_{\alpha,\mathrm{R}}[\ell,k],\mathbf{g}_{\alpha,\mathrm{I}}[\ell,k]]\in\mathbb{C}^{N\times 6}$. Here, the columns fulfill
$\mathbf{g}_{v}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\left(\jmath 2\pi\ell T\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10a)
$\mathbf{g}_{\tau}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\left(-\jmath 2\pi k\Delta f\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10b)
$\mathbf{g}_{\theta_{\mathrm{az}}}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{\mathrm{az}}},$ (10c)
$\mathbf{g}_{\theta_{\mathrm{el}}}\left[\ell,k\right]$ $=\alpha_{s,u}\left(\frac{\partial\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]}{\partial\theta_{s,u}^{\mathrm{el}}}\mathbf{a}\left(\bm{\theta}_{s,u}\right)\right.$ (10d)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left.\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{s,u}^{\mathrm{el}}}\right),$
$\mathbf{g}_{\alpha,\mathrm{R}}\left[\ell,k\right]$ $=\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10e)
$\mathbf{g}_{\alpha,\mathrm{I}}\left[\ell,k\right]$ $=\jmath\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right).$ (10f)
Based on the signal received from the $s$-th satellite, the FIM of ${\bm{\eta}}_{s,u}$ can be computed by the Slepian-Bangs formula as
$\mathbf{J}_{s}({\bm{\eta}}_{s,u})$
$=$ $\frac{2}{N_{0}\Delta f}\sum_{\ell=1}^{L_{\text{p}}}\sum_{k=1}^{K}\Re\Big{\{}\Big{(}\frac{\partial{\mu}_{s,u}\left[\ell,k\right]}{\partial{\bm{\eta}}_{s,u}}\Big{)}\Big{(}\frac{\partial{\mu}_{s,u}\left[\ell,k\right]}{\partial{\bm{\eta}}_{s,u}}\Big{)}^{\mathsf{H}}\Big{\}},$ (11)
$=$ $\frac{2}{N_{0}\Delta f}\sum_{\ell=1}^{L_{\text{p}}}\sum_{k=1}^{K}\Re\big{\{}\mathbf{G}_{s,u}^{\mathsf{T}}\left[\ell,k\right]\mathbf{f}_{s}\left[\ell,k\right]\mathbf{f}_{s}\left[\ell,k\right]^{\mathsf{H}}\mathbf{G}_{s,u}^{*}\left[\ell,k\right]\big{\}}.$
III-B3 The FIM of $\acute{\bm{\eta}}_{u}$
Partitioning $\mathbf{J}_{s}(\bm{\eta}_{s,u})=[\mathbf{X},\mathbf{Y};\mathbf{Y}^{\mathsf{T}},\mathbf{Z}]$, where $\mathbf{X}\in\mathbb{R}^{2\times 2}$, we can compute the FIM of the total utilized channel-domain parameters $\acute{\bm{\eta}}_{s,u}$ as $\mathbf{J}_{s}\left(\acute{\bm{\eta}}_{s,u}\right)=\mathbf{X}-\mathbf{Y}\mathbf{Z}^{-1}\mathbf{Y}^{\mathsf{T}}$. Then, we can obtain the FIM of $\acute{\bm{\eta}}_{u}$ as
$\mathbf{J}\left(\acute{\bm{\eta}}_{u}\right)=\text{blkdiag}\big{\{}\mathbf{J}_{1}\left(\acute{\bm{\eta}}_{1,u}\right),\ldots,\mathbf{J}_{S}\left(\acute{\bm{\eta}}_{S,u}\right)\big{\}}.$ (12)
Based on this, we can infer that the achievable lowest estimation error on $\acute{\bm{\eta}}_{u}$ regardless of the adopted algorithms is characterized by the Gaussian distribution $\mathcal{N}\big{(}\mathbf{0},\mathbf{J}_{s}^{-1}(\acute{\bm{\eta}}_{u})\big{)}$. Any algorithm achieving this error covariance matrix is called an efficient estimator.
III-B4 MCRB and LB Derivation
Suppose an efficient estimator is utilized to estimate $\acute{\bm{\eta}}_{u}$. We now focus on estimating the position-domain parameters, where the aforementioned model mismatches occur. We consider a scenario where each satellite-UT pair has a distinct clock bias $b_{s,u}$ and CFO $\delta_{s,u}$. For convenience, we use $b_{u}$ and $\delta_{u}$ to denote an average clock bias and CFO among the $S$ satellites, thus we can rewrite these individual clock biases and CFOs as $b_{s,u}=b_{u}+\Delta b_{s,u}$ and $\delta_{s,u}=\delta_{u}+\Delta\delta_{s,u}$, respectively, where $\Delta b_{s,u}$ and $\Delta\delta_{s,u}$ denote the differences on each clock bias and CFO.
Let $\mathbf{v}_{s}\in\mathbb{R}^{3}$ be the velocity of the $s$-th LEO satellite. Here, as the velocity of LEO satellite is usually much higher than that of the UT, we ignore the velocity contribution from the UT side. Based on the underlying geometric relationship, the forward model between the entries of $\acute{\bm{\eta}}_{u}$ and unknown position of the $u$-th user can be expressed as
$\upsilon_{s,u}$ $=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u}+\Delta\delta_{s,u},$ (13a)
$\tau_{s,u}$ $=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}+\Delta b_{s,u},$ (13b)
which we refer to as the true model. In real applications, we often ignore these clock bias and CFO differences, i.e., assuming all satellite-UT pairs have the same clock bias and CFO, for algorithmic simplicity. To account for the impact of such model mismatch, here we derive the MCRB. When neglecting the difference on each clock bias and CFO, the mismatched model is given by
$\upsilon_{s,u}$ $=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u},$ (14a)
$\tau_{s,u}$ $=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}.$ (14b)
Now, we suppose a mismatched estimator that aims to estimate the unknown position of the $u$-th user based on the realistic observation $\acute{\bm{\eta}}_{u}$ and simplified model (14) . the total unknown parameters in the mismatch model can be concatenated as $\mathbf{r}_{u}=[\mathbf{p}_{u}^{\mathsf{T}},\delta_{u},b_{u}]^{\mathsf{T}}\in\mathbb{R}^{5}$. Based on the true model (13) and mismatched model (14), we can respectively express the true likelihood function $\xi_{\mathrm{T}}$ and the mismatched likelihood function $\xi_{\mathrm{M}}$ as
$\ln{\xi_{\mathrm{T}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto\big{(}\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)}^{\mathsf{T}}\bm{\Sigma}^{-1}\big{(}\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)},$ (15)
and
$\ln{\xi_{\mathrm{M}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto\big{(}\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)}^{\mathsf{T}}\bm{\Sigma}^{-1}\big{(}\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)},$ (16)
respectively, where the function $\bar{\mathbf{f}}(\cdot)$ maps $\mathbf{r}_{u}$ to $\acute{\bm{\eta}}_{u}$ according to (13), and the function $\tilde{\mathbf{f}}(\cdot)$ maps $\mathbf{r}_{u}$ to $\acute{\bm{\eta}}_{u}$ according to (14). As mentioned in Section III-B3, $\acute{\bm{\eta}}_{u}$ is the estimated channel-domain parameters through an efficient estimator, i.e., $\acute{\bm{\eta}}_{u}\sim\mathcal{N}(\bar{\mathbf{f}}(\mathbf{r}_{u}),\bm{\Sigma})$, where $\bm{\Sigma}=\mathbf{J}^{-1}(\acute{\bm{\eta}}_{u})$.
  • Pseudo-True Parameter: According to (6), the pseudo-true parameter, denoted by $\tilde{\mathbf{r}}_{u}$, is given by
    $\tilde{\mathbf{r}}_{u}$ $=\arg\min_{\mathbf{r}_{u}}\ \mathcal{D}\big{\{}{\xi_{\mathrm{T}}}(\acute{\bm{\eta}}_{u};\bar{\mathbf{r}}_{u})\|{\xi_{\mathrm{M}}}(\acute{\bm{\eta}}_{u};\mathbf{r}_{u})\big{\}},$ (17)
    $=\arg\min_{\mathbf{r}_{u}}\ \big{(}\bar{\mathbf{f}}(\bar{\mathbf{r}}_{u})-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)}^{\mathsf{T}}\bm{\Sigma}^{-1}\big{(}\bar{\mathbf{f}}(\bar{\mathbf{r}}_{u})-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)},$
    where the top bar is used to highlight the ground-truth position-domain parameter and forward model. The above problem can be solved using gradient descent with backtracking line search, wherein the true parameters serve as the initial point[35].
  • MCRB: According to (III-A) and (III-A), the generalized FIMs, i.e., $\mathbf{A}_{\tilde{\mathbf{r}}_{u}}$ and $\mathbf{B}_{\tilde{\mathbf{r}}_{u}}$, can be computed by determining the first- and second-order derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to the elements of $\mathbf{r}_{u}$, as detailed in Appendix A. Subsequently, the MCRB is expressed as
    $\text{MCRB}\left(\tilde{\mathbf{r}}_{u}\right)=\mathbf{A}_{\tilde{\mathbf{r}}_{u}}^{-1}\mathbf{B}_{\tilde{\mathbf{r}}_{u}}\mathbf{A}_{\tilde{\mathbf{r}}_{u}}^{-1}.$ (18)
  • LB: Finally, the estimation error LB matrix can be computed according to (5) as
    $\text{LBM}\left(\tilde{\mathbf{r}}_{u}\right)=\text{MCRB}\left(\tilde{\mathbf{r}}_{u}\right)+\text{Bias}\left(\tilde{\mathbf{r}}_{u}\right),$ (19)
    where $\text{Bias}\left(\tilde{\mathbf{r}}_{u}\right)=\left(\bar{\mathbf{r}}_{u}-\tilde{\mathbf{r}}_{u}\right)\left(\bar{\mathbf{r}}_{u}-\tilde{\mathbf{r}}_{u}\right)^{\mathsf{T}}$. Based on (19), the LB for the expected root mean squared error (RMSE) of the $u$-th UT’s position estimation in the presence of model mismatch is given by
    $\text{LB}\left(\mathbf{p}_{u}\right)=\sqrt{\text{Tr}\big{(}\left[\text{LBM}\left(\tilde{\mathbf{r}}_{u}\right)\right]_{1:3,1:3}\big{)}}.$ (20)
III-C Small-Timescale Uplink Channel Estimation
III-C1 Signal Model
After obtaining the positions of UTs, these large-timescale parameters are fed back to the LEO satellites. The second phase of PACE then focuses on estimating the small-timescale parameters (i.e., the channel gains) at the LEO satellites through uplink channel estimation, followed by complete channel reconstruction.
Let $\mathbf{d}_{s,u}[\ell,k]=\mathbf{a}(\bm{\theta}_{s,u})\sqrt{P_{u}/K}t_{u}\left[\ell,k\right]$, where $P_{u}$ denotes the transmit power of the $u$-th UT, and $t_{u}[\ell,k]$ represents the unit-modulus pilot transmitted by the $u$-th UT during the $\ell$-th symbol and the $k$-th subcarrier. Uniform power allocation is employed across all subcarriers. As previously mentioned, the Doppler shift can be largely mitigated in LEO satellite communications. After being exploited in downlink positioning, it is excluded from the signal model including uplink channel estimation and downlink communication for simplicity. Here, we also assume UTs are multiplexed orthogonally using TDMA. By channel reciprocity, the signal received by the $s$-th LEO satellite from the $u$-th UT is expressed as
$\mathbf{y}_{s,u}\left[\ell,k\right]=\underbrace{\vartheta_{s,u}e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\mathbf{d}_{s,u}[\ell,k]}_{\bar{\mathbf{g}}_{s,u}\left[\ell,k\right]}+\mathbf{z}_{s}\left[\ell,k\right],$ (21)
where $\vartheta_{s,u}=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})$ and $\mathbf{z}_{s}[\ell,k]\sim\mathcal{CN}(\mathbf{0},N_{0}\Delta f\mathbf{I}_{N})$ represents the AWGN.
We observe that $\mathbf{d}_{s,u}[\ell,k]$ is directly determined by the UTs’ position feedback. Consequently, instead of estimating the full channel vector, it suffices to estimate only two scalars, i.e., the channel gain $\vartheta_{s,u}$ and the TOA $\tau_{s,u}$. This significantly simplifies the uplink channel estimation task. However, since the position information is based on estimation rather than ground truth, model mismatch arises again. The received signal based on the estimated position information (mismatched model) is expressed as
$\mathbf{y}_{s,u}\left[\ell,k\right]=\underbrace{\vartheta_{s,u}e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\tilde{\mathbf{d}}_{s,u}\left[\ell,k\right]}_{\tilde{\mathbf{g}}_{s,u}\left[\ell,k\right]}+\mathbf{z}_{s}\left[\ell,k\right],$ (22)
where $\tilde{\mathbf{d}}_{s,u}[\ell,k]=\mathbf{a}(\tilde{\bm{\theta}}_{s,u})\sqrt{P_{u}/K}t_{u}\left[\ell,k\right]$. Here, $\tilde{\bm{\theta}}_{s,u}$ represents the calculated angle-of-arrival (AOA) based on the estimated $\tilde{\mathbf{p}}_{u}$ and the geometrical relationship, given by
[…]
$\tilde{\theta}_{s,u}^{\mathrm{el}}$ $=\text{asin}\Big{(}\frac{\left[\left(\tilde{\mathbf{p}}_{u}-\mathbf{p}_{s}\right)\right]_{3}}{\left\|\tilde{\mathbf{p}}_{u}-\mathbf{p}_{s}\right\|}\Big{)}.$ (23b)
III-C2 MCRB and LB Derivation
We denote the total temporal and spectral observation as $\mathbf{y}_{s,u}=[\mathbf{y}_{s,u}^{\mathsf{T}}[1,1],\ldots,\mathbf{y}_{s,u}^{\mathsf{T}}[1,K],\ldots,\mathbf{y}_{s,u}^{\mathsf{T}}[L_{\text{c}},K]]^{\mathsf{T}}\in\mathbb{C}^{L_{\text{c}}KN}$, where $L_{\text{c}}$ represents the number of pilot symbols used for uplink channel estimation. Let $\bm{\eta}_{s,u}=[\Re(\vartheta_{s,u}),\Im(\vartheta_{s,u}),\tau_{s,u}]^{\mathsf{T}}\in\mathbb{R}^{3}$ collect the parameters to be estimated in the uplink channel estimation. The likelihood functions under the true and mismatched models are given by
$\ln{\xi_{\mathrm{T}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto\left\|\mathbf{y}_{s,u}-\bar{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (24)
and
$\ln{\xi_{\mathrm{M}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto\left\|\mathbf{y}_{s,u}-\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (25)
respectively, where we define $\bar{\mathbf{g}}(\bm{\eta}_{s,u})=[\bar{\mathbf{g}}_{s,u}^{\mathsf{T}}[1,1],\ldots,\bar{\mathbf{g}}_{s,u}^{\mathsf{T}}[1,K],\ldots,\bar{\mathbf{g}}_{s,u}^{\mathsf{T}}[L_{\text{c}},K]]^{\mathsf{T}}\in\mathbb{C}^{L_{\text{c}}KN}$ and $\tilde{\mathbf{g}}(\bm{\eta}_{s,u})=[\tilde{\mathbf{g}}_{s,u}^{\mathsf{T}}[1,1],\ldots,\tilde{\mathbf{g}}_{s,u}^{\mathsf{T}}[1,K],\ldots,\tilde{\mathbf{g}}_{s,u}^{\mathsf{T}}[L_{\text{c}},K]]^{\mathsf{T}}\in\mathbb{C}^{L_{\text{c}}KN}$.
  • Pseudo-True Parameter: According to (6), the pseudo-true parameter, denoted by $\tilde{\bm{\eta}}_{s,u}$, is derived as
    $\tilde{\bm{\eta}}_{s,u}$ $=\arg\min_{\bm{\eta}_{s,u}}\ \mathcal{D}\big{\{}{\xi_{\mathrm{T}}}\left(\mathbf{y}_{s,u};\bar{\bm{\eta}}_{s,u}\right)\|{\xi_{\mathrm{M}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\big{\}},$
    $=\arg\min_{\bm{\eta}_{s,u}}\ \left\|\bar{\mathbf{q}}\left(\bar{\bm{\eta}}_{s,u}\right)-\tilde{\mathbf{q}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (26)
    where the bar above the parameter emphasizes the ground-truth value. Once again, by employing gradient descent and treating truth parameter as initial point, the above problem can be efficiently solved [35].
  • MCRB: The generalized FIMs, i.e., $\mathbf{A}_{\tilde{\bm{\eta}}_{s,u}}$ and $\mathbf{B}_{\tilde{\bm{\eta}}_{s,u}}$, are obtained by calculating the first- and second-order derivatives of $\tilde{\mathbf{q}}\left(\bm{\eta}_{s,u}\right)$ with respect to the elements of $\bm{\eta}_{s,u}$, as described in Appendix B. Based on these, the MCRB is given by
    $\text{MCRB}\left(\tilde{\bm{\eta}}_{s,u}\right)=\mathbf{A}_{\tilde{\bm{\eta}}_{s,u}}^{-1}\mathbf{B}_{\tilde{\bm{\eta}}_{s,u}}\mathbf{A}_{\tilde{\bm{\eta}}_{s,u}}^{-1}.$ (27)
  • LB: The estimation error LB matrix is then calculated as
    $\text{LBM}\left(\tilde{\bm{\eta}}_{s,u}\right)=\text{MCRB}\left(\tilde{\bm{\eta}}_{s,u}\right)+\text{Bias}\left(\tilde{\bm{\eta}}_{s,u}\right),$ (28)
    where $\text{Bias}\left(\tilde{\bm{\eta}}_{s,u}\right)=\left(\bar{\bm{\eta}}_{s,u}-\tilde{\bm{\eta}}_{s,u}\right)\left(\bar{\bm{\eta}}_{s,u}-\tilde{\bm{\eta}}_{s,u}\right)^{\mathsf{T}}$. For an efficient estimator, the LB for the expected RMSE of the channel gain $\vartheta_{s,u}$ and the TOA $\tau_{s,u}$, in the presence of model mismatch, are expressed as
    $\text{LB}\left(\vartheta_{s,u}\right)=\sqrt{\text{Tr}\big{(}\left[\text{LBM}\left(\tilde{\bm{\eta}}_{s,u}\right)\right]_{1:2,1:2}\big{)}},$ (29)
    and
    $\text{LB}\left(\tau_{s,u}\right)=\sqrt{\left[\text{LBM}\left(\tilde{\bm{\eta}}_{s,u}\right)\right]_{3,3}},$ (30)
    respectively.
IV Distributed Beamforming for Downlink Communications
The proposed two-timescale PACE enables channel reconstruction between each LEO satellite and all UTs, enabling CSI-based beamforming for downlink communication. To address the limited link budget in single-satellite downlink scenarios, we introduce distributed multi-LEO satellite beamforming where each UT is served by the same satellite cluster. This distributed approach eliminates the need for inter-satellite CSI sharing, as beamforming decisions are made locally at each satellite using only its CSI, rather than requiring consolidated global CSI, which significantly reduces system complexity and latency.
IV-A Signal Model
Without centralized coordination, we manage inter-satellite interference by allocating orthogonal frequency bands to different LEO satellites. The total subcarrier set $\Phi=\{1,2,\ldots,K\}$ is partitioned into $S$ equal subsets $\Phi_{1},\Phi_{2},\ldots,\Phi_{S}$ corresponding to $S$ LEO satellites. For each satellite $s$, its subset is defined as $\Phi_{s}=\{(s-1)\bar{K}+1,(s-1)\bar{K}+1,\ldots,s\bar{K}\}$, where $\bar{K}=K/S$. For the $s$-th LEO satellite transmitting over the $k$-th subcarrier ($k\in\Phi_{s}$), the signal is given by
$\mathbf{x}_{s}\left[k\right]=\mathbf{W}_{s}\mathbf{s}\left[k\right],$ (31)
where $\mathbf{W}_{s}=[\mathbf{w}_{s,1},\mathbf{w}_{s,2},\ldots,\mathbf{w}_{s,U}]\in\mathbb{C}^{N\times U}$ represents the beamformer [^3] We adopt subcarrier-coherent beamforming, where the beamformer is consistent across subcarriers, with only amplitude adjustments allowed. This approach strikes a balance between performance and computational efficiency, which is critical given the limited onboard processing capabilities of LEO satellites. Although less flexible than subcarrier-specific beamforming, it substantially reduces optimization complexity in systems with a large number of subcarriers. The framework can be extended to subcarrier-specific beamforming, albeit at the expense of increased computational complexity. for the $s$-th LEO satellite, and $\mathbf{s}\left[k\right]\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{U})$ denotes the data streams transmitted to $U$ UTs over the $k$-th subcarrier.
The received signal at the $u$-th UT from the $s$-th LEO satellite over the $k$-th subcarrier can be represented as
$y_{u}\left[k\right]=$ $\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,u}s_{u}\left[k\right]$
$+\underbrace{\sum_{m=1,m\neq u}^{U}\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,m}s_{m}\left[k\right]}_{\text{Inter-user interference}}+z_{u}\left[k\right],$ (32)
where $\mathbf{h}_{s,u}[k]=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\mathbf{a}(\bm{\theta}_{s,u})$ denotes the channel vector incorporating Doppler shift compensation. The term $z_{u}[k]\sim\mathcal{CN}(\mathbf{0},N_{0}\Delta f)$ represents the AWGN. The achievable communication rate for the $u$-th UT, aggregated across all satellites and assigned subcarriers, is given by
$R_{u}=\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}\left[k\right],$ (33)
where $R_{s,u}\left[k\right]$ represents the communication rate between the $s$-th LEO satellite and the $u$-th UT over the $k$-th subcarrier, calculated as
$R_{s,u}\left[k\right]=\log_{2}\left(1+\frac{\left|\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,u}\right|^{2}}{I_{s,u}\left[k\right]+N_{0}\Delta f}\right),$ (34)
where $I_{s,u}[k]=\sum_{m=1,m\neq u}^{U}|\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,m}|^{2}$.
IV-B Problem Formulation
We aim to optimize the beamformers at each LEO satellite to maximize the sum rate of all UTs. The optimization problem is formulated as follows
$\mathop{\max}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}R_{u}=\sum_{u=1}^{U}\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}[k]$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}},\;\forall s,$ (35)
where $P_{s}$ denotes the power budget at the $s$-th LEO satellite.
Note that, because $\sum_{u=1}^{U}\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}[k]=\sum_{s=1}^{S}\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}R_{s,u}[k]$, the problem in (IV-B) is inherently distributed. Consequently, a subproblem corresponding to the $s$-th LEO satellite can be expressed as
$\mathop{\max}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}R_{s,u}[k]$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}}.$ (36)
By solving (IV-B) for each LEO satellite, the original problem in (IV-B) can be effectively addressed.
IV-C Proposed Distributed Beamforming Approach
In the following, we employ the weighted sum mean squared error minimization (WMMSE) optimization framework [36] to solve (IV-B) iteratively. Specifically, by introducing the auxiliary variables $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$, where $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$ serve as the MMSE equalizer and weight, respectively [36], (IV-B) can be equivalently reformulated as
$\mathop{\max}\limits_{\mu_{s,u}[k],\omega_{s,u}[k],\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}\log_{2}\omega_{s,u}[k]-\omega_{s,u}[k]\Psi_{s,u}[k]$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}},$ (37)
where $\Psi_{s,u}[k]=|1-\mu_{s,u}[k]\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,u}|^{2}+|\mu_{s,u}[k]|^{2}(I_{s,u}[k]+N_{0}\Delta f)$. The optimization problem in (IV-C) is solved by iteratively updating $\mathbf{W}_{s}$, $\omega_{s,u}[k]$, and $\mu_{s,u}[k]$, as detailed below.
IV-C1 Update $\mu_{s,u}[k]$
With $\omega_{s,u}[k]$ and $\mathbf{W}_{s}$ fixed, the optimal $\mu_{s,u}[k]$ is derived by setting $\partial\Psi_{s,u}[k]/\partial\mu_{s,u}[k]=0$, resulting in
$\mu_{s,u}[k]=\frac{\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,u}}{I_{s,u}[k]+N_{0}\Delta f}.$ (38)
IV-C2 Update $\omega_{s,u}[k]$
With $\mu_{s,u}[k]$ and $\mathbf{W}_{s}$ fixed, the optimal $\omega_{s,u}[k]$ is obtained by setting $\partial\Psi_{s,u}[k]/\partial\omega_{s,u}[k]=0$, leading to
$\omega_{s,u}[k]=\frac{1}{\Psi_{s,u}[k]}.$ (39)
IV-C3 Update $\mathbf{W}_{s}$
With $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$ fixed, the optimization problem to find the optimal $\mathbf{W}_{s}$ is formulated as
$\mathop{\min}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}\Psi_{s,u}[k]$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}}.$ (40)
Here, $\Psi_{s,u}[k]$ can be expressed as the sum of a positive semi-definite quadratic form and a linear term. Consequently, (IV-C3) is a convex quadratically constrained quadratic program (QCQP) problem with respect to $\mathbf{W}_{s}$, which can be efficiently solved via convex optimization toolboxes such as CVX.
The process of solving (IV-B) is outlined in Algorithm 1. The initial beamformer $\mathbf{W}_{s}$ for the $s$-th LEO satellite can be constructed using maximum ratio transmission (MRT), which aligns the beamformer with the reconstructed channels (by excluding the frequency-dependent delay-induced phase rotation). Specifically, we have $\mathbf{W}_{s}=\zeta_{s}[\hat{\vartheta}_{s,1}^{*}\mathbf{a}^{*}(\hat{\bm{\theta}}_{s,1}),\ldots,\hat{\vartheta}_{s,U}^{*}\mathbf{a}^{*}(\hat{\bm{\theta}}_{s,U})]$, where the hat notation indicates estimated values and $\zeta_{s}$ is a normalization factor that ensures the fulfillment of the power budget. Since the WMMSE optimization framework is utilized to solve each subproblem, i.e., (IV-B), its convergence is well-established [36]. The computational complexity of each iteration is primarily determined by solving (IV-C3), which, due to its QCQP nature, is characterized by $\mathcal{O}(N^{3}U^{3})$.
Algorithm 1 Algorithm for Solving (IV-B)
1:Initialize: $\mathbf{W}_{s},\forall s$;
2:for $s=1:S$ do
3:     repeat
4:         Update $\mu_{s,u}[k]$ using (38);
5:         Update $\omega_{s,u}[k]$ using (39);
6:         Update $\mathbf{W}_{s}$ by solving (IV-C3) via CVX;
7:     until the reduction ratio of the objective value falls below a predefined threshold;
8:end for
9:Output: $\mathbf{W}_{s},\forall s$.
TABLE I: Simulation Parameters
Parameter Value
[…]
Subcarrier number $K$ 1024
[…]
Number of LEO satellites $S$ 8
Number of UTs $U$ 4
Antenna number $N=N_{\text{h}}\times N_{\text{v}}$ $8\times 8$
Antenna radiation gain $G(\cdot)$ $\cos(\cdot)$
V Numerical Results
V-A Simulation Parameters
Figure 4: Illustrative example of the considered scenario.
[…]
V-B Compared Schemes
V-B1 Downlink Positioning
Since we do not optimize the beamforming matrix $\mathbf{F}_{s}$, it is generated directly based on two principles, which differ depending on whether position information is utilized:
  • Position-Aided Beamforming (PAB): Leveraging the position information of UTs obtained from the previous position-coherent frame, each LEO satellite constructs $\mathbf{F}_{s}$ as a matrix whose columns correspond to the conjugates of the steering vectors for the UTs’ estimated directions. Specifically, $\mathbf{F}_{s}=\bar{\zeta}_{s}[\mathbf{a}^{*}(\hat{\bm{\theta}}_{s,1}),\ldots,\mathbf{a}^{*}(\hat{\bm{\theta}}_{s,U})]$, where $\bar{\zeta}_{s}$ is a normalization factor that ensures compliance with the power budget.
  • Vertically Directed Beamforming (VDB): In the absence of position information, each LEO satellite generates a vertical beam directed toward the Earth’s core. In this case, the beamforming matrix reduces to a steering vector with an elevation angle of $\pi/2$.
V-B2 Uplink Channel Estimation
Depending on the availability of position information at the LEO satellites, we consider three types of schemes for evaluating the performance of uplink channel estimation:
  • PACE: This is the proposed two-timescale hybrid downlink positioning and uplink channel estimation scheme. Its performance is rigorously characterized through MCRB analysis.
  • Perfect-PACE: This represents an ideal benchmark where perfect position information of the UTs is assumed to be available at the LEO satellites. In this case, the performance of uplink channel estimation is evaluated without model mismatch and analyzed using the CRB.
  • Uplink Channel Estimation (UCE): As previously discussed, in the absence of position information, the channel vector, i.e., $\mathbf{h}_{s,u}[k]=e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})\mathbf{a}(\bm{\theta}_{s,u})$, can be directly estimated in the uplink using a geometry-agnostic approach. By defining $\bm{h}_{s,u}=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})\mathbf{a}(\bm{\theta}_{s,u})$, we analyze the performance limits of estimating $\bm{h}_{s,u}$ and $\tau_{s,u}$ simultaneously by deriving the CRB. This derivation extends the approach in [40], which was originally developed for a single-carrier system.
V-B3 Downlink Communication
To evaluate the performance of downlink communication, we utilize the CSI obtained through the different channel estimation schemes introduced above to perform distributed beamforming, as summarized in Algorithm 1. The resulting beamformer is tested using ground-truth CSI, yielding the actual sum rate. Additionally, we consider two alternative benchmarks:
  • PAB: This approach is identical to the one introduced for downlink positioning, which relies solely on position information. However, the beamformer at each LEO satellite is applied only to a subset of subcarriers.
  • Single-Satellite Service ($\text{S}^{3}$): In this method, each UT is served by only a single LEO satellite. Specifically, the one providing the strongest signal strength, rather than being served by all satellites within a cluster.
V-C Downlink Positioning
$10^{-1}$$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$er-LEO satellite power budget [dBm]RMSE [m]PAB: LBPAB: MCRBPAB: BiasPAB: CRBVDB: LBVDB: MCRBVDB: BiasVDB: CRB
Figure 5: Positioning RMSE LB versus per-LEO satellite power budget.
As shown in Fig. 5, we compare the positioning error bounds and biases under PAB and VDB as functions of the per-satellite power budget. The results reveal that the LB of positioning error without model mismatch (represented by the positioning CRB under perfectly time- and frequency-synchronized satellites) decreases with increasing power. In contrast, the LB with mismatch initially decreases but eventually saturates at a value determined by the bias, which remains independent of the transmit power. This emphasizes the detrimental impact of imperfect time- and frequency-synchronization among LEO satellites on positioning performance [26, 24], potentially leading to overly optimistic performance estimates. Moreover, compared to VDB, PAB achieves a lower LB, with the performance gain being more pronounced at lower (and more practical) transmit power levels. This advantage arises because PAB, by utilizing position information, can allocate more power to the UTs than VDB, underscoring the importance of integrating prior position knowledge into the positioning system.
[…]
Figure 6: Positioning RMSE LB versus maximum clock bias and CFO.
As illustrated in Fig. 6, we further analyze the impact of varying time and frequency biases among satellite-UT pairs on positioning performance, resulting from imperfect time and frequency synchronization among LEO satellites. Specifically, we present the LB of positioning error (under PAB) as a function of the maximum clock bias and CFO values. The clock bias and CFO for each satellite-UT pair are uniformly distributed between 00 and their respective maximum values. The results reveal that positioning performance deteriorates significantly as the maximum clock bias and CFO increase. A slight clock bias (on the order of milliseconds) or a minor CFO (on the order of $\Delta f/10$) can result in positioning errors of several kilometers, highlighting the critical importance of maintaining precise synchronization among LEO satellites in both time and frequency domains to achieve high-accuracy positioning.
$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{-15}$$10^{-13}$$10^{-11}$$10^{-9}$osition error [m]RMSECRBLBMCRBBias
(a) Channel gain estimation
$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{-13}$$10^{-11}$$10^{-9}$osition error [m]RMSE [s]
(b) TOA estimation
Figure 7: Uplink channel estimation RMSE LB versus positioning error: (a) Channel vector estimation; (b) TOA estimation.
[…]
(a) Channel vector estimation
[…]
(b) TOA estimation
Figure 8: Uplink channel estimation RMSE LB versus per-UT power budget: (a) Channel vector estimation; (b) TOA estimation.
V-D Uplink Channel Estimation
In Figs. 7(a) and (b), we evaluate the LB for both channel gain and TOA estimation RMSE using our proposed PACE across varying positioning errors, comparing it with Perfect-PACE (without positioning error-induced model mismatch). The results reveal that when positioning errors are small, PACE’s LB closely matches that of Perfect-PACE for both parameters. However, as positioning errors increase, PACE’s performance deteriorates due to growing model mismatch. For channel gain estimation, we observe that while the MCRB-induced error remains relatively insensitive to positioning errors, the bias component increases proportionally with positioning error, eventually becoming the primary contributor to overall channel gain error. The TOA estimation exhibits different behavior: its bias initially increases with positioning error before reaching a plateau, while the MCRB-induced error continues to grow, ultimately becoming the dominant factor in the total TOA estimation error. These findings underscore the importance of accurate downlink positioning in the first stage of PACE for achieving high-accuracy uplink channel estimation in the second stage.
In Figs. 8(a) and (b), we compare the LB for the RMSE of both channel vector and TOA estimation using the proposed PACE against those obtained under UCE, versus UT’s power budget. Note that PACE estimates both the channel gain $\alpha_{s,u}$ and the TOA $\tau_{s,u}$, whereas UCE estimates the channel vector $\bm{h}_{s,u}=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})\mathbf{a}(\bm{\theta}_{s,u})$ (excluding the TOA-induced phase) in addition to $\tau_{s,u}$. To enable a unified comparison, we transform the LB for channel gain estimation RMSE in PACE into the corresponding LB for channel vector estimation RMSE (using the form $\hat{\bm{h}}_{s,u}=\hat{\alpha}_{s,u}G(\hat{\theta}_{s,u}^{\mathrm{el}})\mathbf{a}(\hat{\bm{\theta}}_{s,u})$) through algebraic manipulation. As shown, the TOA estimation performance of PACE closely aligns with that of UCE, exhibiting only negligible degradation due to positioning error. This alignment occurs because TOA estimation is independent of whether the other unknown parameter is the channel gain or the channel vector. This independence can be readily demonstrated by deriving the CRB for TOA estimation under Perfect-PACE and UCE, which we omit here due to space limitations. In contrast, PACE significantly outperforms UCE in channel vector estimation. This superiority is attributed to PACE’s utilization of position information, which reduces the complexity of estimating a vector (as required by UCE) to that of estimating a scalar, albeit subject to positioning error. Consequently, PACE achieves higher channel estimation accuracy under the same pilot length and power budget.
[…]
Figure 9: Sum rate versus per-LEO satellite antenna number.
[…]
Figure 10: Sum rate versus per-UT power budget.
V-E Donwlink Communication
In Fig. 9, we compare the downlink communication sum rates under various approaches versus the number of antennas equipped on each LEO satellite. As observed, the sum rate under the proposed PACE significantly outperforms those under UCE and other baselines, reaching a level very close to the sum rate with perfect CSI. This results from PACE’s superior performance in reconstructing the channel accurately. Meanwhile, the sum rate under PAB is even lower than that under $\text{S}^{3}$, underscoring the importance of accurate CSI in delivering the full capability of multi-LEO satellite downlink communication. Furthermore, with an increasing number of antennas on each LEO satellite, the performance gain of the proposed scheme becomes more evident, highlighting its potential in facilitating high-speed LEO satellite communication with massive antenna arrays.
As shown in Fig. 10, we compare the resultant sum rates under various schemes versus the per-UT power budget. It can be seen that the sum rates under PAB and $\text{S}^{3}$ are irrelevant to UTs’ transmit power, while $\text{S}^{3}$ remains superior to PAB throughout the entire range. However, with increasing transmit power from the UT side, the sum rates under PACE and UCE increase and gradually approach those under perfect CSI. This is because the channel estimation accuracy under both schemes increases with higher transmit power, which translates into a sum rate gain. Nevertheless, the sum rate under PACE is significantly always higher than that under UCE (especially notable when the transmit power at the UT is low). This highlights the importance of leveraging position information to obtain accurate CSI when the power budget at the UTs is limited, which is particularly relevant for equipments such as cell phones.
VI Conclusion
This paper presents a novel framework for integrating positioning and communication in multi-LEO satellite systems, leveraging positioning capabilities to enhance channel estimation and thereby facilitate communication. By capitalizing on the distinct timescale variations of position-related parameters and channel gains, we propose a two-timescale frame structure that strategically schedules downlink positioning, uplink channel estimation, and downlink communication. A rigorous MCRB-based analysis characterizes the impact of practical imperfections, including inter-satellite clock bias and CFO, on positioning accuracy. Additionally, we theoretically demonstrate through MCRB analysis how position information improves uplink channel estimation even under inevitable positioning errors. To overcome the limited link budget offered by single-satellite-based downlink communication, we introduce a multi-satellite distributed beamforming strategy that relies on local CSI at each satellite, eliminating the need for centralized coordination. Numerical results validate our theoretical findings. As a direction for future work, developing practical algorithms and validating the framework through experimental testing would be valuable next steps.
Appendix A Derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to $\mathbf{r}_{u}$
1) Non-Zero First-Order Derivative:
$\frac{\partial\upsilon_{s,u}}{\partial\mathbf{p}_{u}}$ $=\frac{\mathbf{v}_{s}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}-\frac{\left(\mathbf{v}_{s}^{\mathsf{T}}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\right)\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}},$
$\frac{\partial\tau_{s,u}}{\partial\mathbf{p}_{u}}$ $=\frac{\mathbf{p}_{u}-\mathbf{p}_{s}}{c\|\mathbf{p}_{u}-\mathbf{p}_{s}\|},\quad\frac{\partial\upsilon_{s,u}}{\partial\delta_{u}}=1,\quad\frac{\partial\tau_{s,u}}{\partial b_{u}}=1.$
2) Non-Zero Second-Order Derivative:
$\frac{\partial^{2}\upsilon_{s,u}}{\partial\mathbf{p}_{u}^{2}}$ $=-\frac{\mathbf{v}_{s}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}}$
$-\frac{\mathbf{v}_{s}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}+\mathbf{v}_{s}^{\mathsf{T}}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\mathbf{I}_{3}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}},$
$+\frac{3\left(\mathbf{v}_{s}^{\mathsf{T}}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\right)\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{5}},$
$\frac{\partial^{2}\tau_{s,u}}{\partial\mathbf{p}_{u}^{2}}$ $=\frac{\mathbf{I}_{3}}{c\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}-\frac{\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}}{c\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}}.$
Appendix B Derivatives of $\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)$ with respect to $\bm{\eta}_{s,u}$
Note that the derivatives $\partial\tilde{\mathbf{g}}(\bm{\eta}_{s,u})/\partial[\bm{\eta}_{s,u}]_{i}$ and $\partial^{2}\tilde{\mathbf{g}}(\bm{\eta}_{s,u})/(\partial[\bm{\eta}_{s,u}]_{i}\partial[\bm{\eta}_{s,u}]_{j})$ can be constructed by concatenating $\partial\tilde{\mathbf{g}}_{s,u}[\ell,k]/\partial[\bm{\eta}_{s,u}]_{i}$ and $\partial^{2}\tilde{\mathbf{g}}_{s,u}[\ell,k]/(\partial[\bm{\eta}_{s,u}]_{i}\partial[\bm{\eta}_{s,u}]_{j})$, respectively, for $\ell=1,\dots,L_{\text{c}}$ and $k=1,\dots,K$. For simplicity, we provide the latter below.
1) Non-Zero First-Order Derivative:
$\frac{\partial\tilde{\mathbf{g}}_{s,u}\left[\ell,k\right]}{\partial\Re\left\{\vartheta{s,u}\right\}}$ $=e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\tilde{\mathbf{d}}_{s,u}\left[\ell,k\right],$
$\frac{\partial\tilde{\mathbf{g}}_{s,u}\left[\ell,k\right]}{\partial\Im\left\{\vartheta_{s,u}\right\}}$ $=\jmath e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\tilde{\mathbf{d}}_{s,u}\left[\ell,k\right],$
$\frac{\partial\tilde{\mathbf{g}}_{s,u}\left[\ell,k\right]}{\partial\tau_{s,u}}$ $=-\vartheta_{s,u}\jmath 2\pi k\Delta fe^{-\jmath 2\pi k\Delta f\tau_{s,u}}\tilde{\mathbf{d}}_{s,u}\left[\ell,k\right].$
2) Non-Zero Second-Order Derivative:
$\frac{\partial^{2}\tilde{\mathbf{g}}_{s,u}\left[\ell,k\right]}{\partial\Re\left\{\vartheta_{s,u}\right\}\partial\tau_{s,u}}$ $=-\jmath 2\pi k\Delta fe^{-\jmath 2\pi k\Delta f\tau_{s,u}}\tilde{\mathbf{d}}_{s,u}\left[\ell,k\right],$
$\frac{\partial^{2}\tilde{\mathbf{g}}_{s,u}\left[\ell,k\right]}{\partial\Im\left\{\vartheta_{s,u}\right\}\partial\tau_{s,u}}$ $=2\pi k\Delta fe^{-\jmath 2\pi k\Delta f\tau_{s,u}}\tilde{\mathbf{d}}_{s,u}\left[\ell,k\right],$
$\frac{\partial^{2}\tilde{\mathbf{g}}_{s,u}\left[\ell,k\right]}{\partial\tau_{s,u}^{2}}$ $=-\vartheta_{s,u}(2\pi k\Delta f)^{2}e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\tilde{\mathbf{d}}_{s,u}\left[\ell,k\right].$

### v4

@@
-Positioning-Aided Channel Estimation for Multi-LEO Satellite Downlink Communications
+Positioning-Aided Channel Estimation for Multi-LEO Satellite Cooperative Beamforming
 Abstract
 We investigate a multi-low Earth orbit (LEO) satellite system that simultaneously provides positioning and communication services to terrestrial user terminals.
-To address the challenges of channel estimation in LEO satellite systems, we propose a novel two-timescale positioning-aided channel estimation framework, exploiting the distinct variation rates of position-related parameters and channel gains inherent in LEO satellite channels.
+To address the challenges of accurately acquiring channel state information in LEO satellite systems, we propose a novel two-timescale positioning-aided channel estimation framework, exploiting the distinct variation rates of position-related parameters and channel gains inherent in LEO satellite channels.
 Using the misspecified Cramér-Rao bound (MCRB) theory, we systematically analyze positioning performance under practical imperfections, such as inter-satellite clock bias and carrier frequency offset.
 Furthermore, we theoretically demonstrate how position information derived from downlink positioning can enhance uplink channel estimation accuracy, even in the presence of positioning errors, through an MCRB-based analysis.
-To overcome the constraints of limited link budgets and communication rates associated with single-satellite-based communication, we develop a distributed beamforming strategy for downlink communication.
-This strategy allows LEO satellites to independently optimize their beamformers using local channel state information, eliminating the need for centralized processing while preserving the advantages of multi-satellite cooperative communication.
-Theoretical analyses and numerical results confirm the effectiveness of the proposed framework in achieving high-precision downlink positioning under practical imperfections, facilitating uplink channel estimation, and enabling efficient downlink communication.
+To address the limited link budgets and communication rates of single-satellite communication, we develop a multi-LEO cooperative beamforming strategy for downlink transmission that leverages cluster-wise satellite cooperation while maintaining reduced complexity.
+Theoretical analyses and numerical results confirm the effectiveness of the proposed framework in facilitating high-precision downlink positioning under practical imperfections, facilitating uplink channel estimation, and enabling efficient downlink communication.
 Index Terms:
-LEO satellite positioning and communication, channel estimation, MCRB, distributed beamforming.
+LEO satellite positioning and communication, channel estimation, MCRB, cooperative beamforming.
 I Introduction
@@
 In these underdeveloped areas, reliable communication and internet access often remain either unavailable or prohibitively expensive, contributing to a growing digital divide[1].
-The non-terrestrial networks (NTN)-based communications have emerged as a promising solution to bridge this gap, with airborne and spaceborne platforms such as high-altitude platform stations and satellites, serving as access points to enable truly global connectivity[2].
-Among these candidates, low Earth orbit (LEO) satellites have attracted particular attention due to their distinct advantages over medium Earth orbit (MEO) and geostationary Earth orbit (GEO) satellites, including reduced path loss and propagation delay, greater constellation flexibility, and lower launch costs[3, 4, 5].
+The non-terrestrial networks (NTN)-based communications have emerged as a promising solution to bridge global connectivity gaps, leveraging airborne and spaceborne platforms, such as high-altitude platforms (HAPs) and satellites, as aerial access points[2, 3, 4, 5, 6, 7].
+Among these platforms, low Earth orbit (LEO) satellites have garnered particular attention due to several advantages over medium Earth orbit (MEO), geostationary Earth orbit (GEO) satellites, and HAPs.
+These advantages include lower path loss and reduced propagation delays, increased flexibility in constellation design, and comparatively lower deployment and launch costs[8, 9, 10].
 The advancement of LEO satellite-based global connectivity presents promising opportunities alongside significant challenges, particularly in channel estimation, a crucial aspect for meeting increasingly demanding high-speed communication requirements[11, 12].
@@
 This extended signal travel distance, combined with limited onboard power resources, significantly constrains the link budget of single-LEO satellite-based communication systems, creating a substantial barrier to achieving the ambitious throughput goals of wideband NTN communications[2, 14, 12].
-Although satellite channels pose significant challenges for CSI acquisition, they exhibit distinct characteristics compared to terrestrial channels, particularly stronger line-of-sight (LOS) dominance and geometric properties [9].
-These characteristics are inherently related to position information [5], suggesting that strategic utilization of UT position information could potentially address channel estimation challenges in LEO satellite communications, especially by enabling partial channel inference through the reconstruction of components such as steering vectors.
-However, this approach raises critical research questions, such as how to achieve accurate and efficient position acquisition within LEO satellite communication networks and how to effectively leverage this position information to enhance LEO satellite channel estimation.
-On the other hand, to address the limited link budget of single-satellite systems, multi-satellite cooperative transmission has emerged as a promising solution [10, 11, 12, 13, 14, 15].
-However, achieving coherent transmission across satellites within non-orthogonal frequency or time resources presents significant challenges, such as phase synchronization among different satellites and the need for centralized processing based on global CSI[10].
-These centralized operations introduce substantial processing overhead and latency, further complicating the practical implementation of cooperative transmission schemes.
+Although satellite channels pose significant challenges for CSI acquisition, they exhibit distinct characteristics compared to terrestrial channels, particularly strong line-of-sight (LOS) dominance and pronounced geometric structures [14].
+These properties are intrinsically tied to position information [10], suggesting that judicious use of UT position knowledge could help alleviate channel estimation challenges in LEO satellite communications, e.g., by enabling partial channel inference through the reconstruction of geometry-dependent components such as steering vectors.
+This viewpoint naturally raises two key questions: (i) how to acquire accurate UT position information efficiently in LEO satellite networks, and (ii) how to leverage such information to improve LEO channel estimation and, ultimately, communication performance.
+On the other hand, to overcome the limited link budget inherent in single-satellite service, multi-satellite cooperative transmission has emerged as a promising solution under diverse design objectives [15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
+Recent works have, for example, designed cooperative transmission schemes to proactively mitigate asynchronous interference in single-carrier/statistical-CSI settings [22], extended such treatments to orthogonal frequency division multiplexing (OFDM) systems (often resorting to high-complexity quadratically constrained quadratic program (QCQP)-based formulations) [23], or emphasized joint power allocation and user scheduling with heuristic precoder constructions [24].
+While these advances are highly relevant, they are not directly aligned with the objective of this paper, which uses cooperative beamforming primarily as a controlled baseline to assess the utility of positioning-aided CSI.
+Moreover, despite the growing interest in multi-satellite cooperation, the computational scalability of cooperative beamforming remains a key bottleneck, especially for resource-constrained LEO satellite platforms.
 The role of position information in terrestrial networks has been widely recognized for enhancing communication performance[25, 26, 27, 28, 29].
@@
 While GNSS appears to be a viable alternative, as demonstrated in [24], where UTs’ positions are obtained via GNSS and fed back to LEO satellites for downlink communication, the critical impact of positioning errors and the reliance on overidealistic synchronization assumptions remains under-explored and warrant further investigation in practical implementations.
-Concurrent with advances in LEO satellite communication, there is growing interest in exploring LEO satellites as potential supplements or alternatives to GNSS [22, 23].
+Concurrent with advances in LEO satellite communication, there is growing interest in exploring LEO satellites as potential supplements or alternatives to GNSS [30, 31, 32, 33, 34].
 This interest arises from the inherent advantages of LEO satellites over GNSS, including stronger signal reception, extensive constellations, and broad frequency diversity, characteristics that have also partially driven the success of LEO satellite communications [31, 35, 36].
@@
 Moreover, it does not account for the inevitable clock bias and carrier frequency offset (CFO) among LEO satellites, factors that may significantly affect positioning performance.
-Three key questions emerge under this context: 1) how can the positioning performance of LEO satellites be effectively characterized under practical imperfections such as clock bias and CFO; 2) how can the position information of UTs obtained through LEO satellite positioning be leveraged to assist the challenging channel estimation task; and 3) how can the benefits of the improved link budget offered by multi-satellite cooperative communication be harnessed while avoiding the complexities of synchronization and centralized processing?
-In this paper, we investigate a scenario where a cluster of LEO satellites, each equipped with a uniform planar array (UPA), simultaneously provides downlink positioning and communication services to multiple single-antenna terrestrial UTs.
-We aim to: 1) characterize the positioning performance under practical impairments such as clock bias/CFO incurred by imperfect time-/frequency-domain synchronization; 2) unveil how the UTs’ position information obtained by downlink positioning can be exploited to enhance uplink channel estimation; and 3) explore distributed beamforming relying solely on local CSI acquired at each satellite, enabling efficient downlink communication while avoiding the overhead and complexity typically associated with centralized operations.
+In this paper, we investigate a cluster of LEO satellites, each equipped with a uniform planar array (UPA), that simultaneously provide downlink positioning and communication services to multiple single-antenna terrestrial UTs.
+We aim to address the key challenges that specifically arise in this setting: characterizing positioning performance under practical synchronization imperfections such as inter-satellite clock bias and CFO; leveraging user position information to enhance channel estimation while accounting for inevitable positioning errors; and mitigating single-satellite link-budget limitations through a carefully designed cooperative multi-satellite beamforming scheme with reduced optimization complexity.
 The main contributions of this paper are as follows.
-  • Building on the distinct variation rates of position-related parameters and random channel gains in LEO satellite channels, as detailed in Section II, we propose a two-timescale frame structure.
+  • Building on the distinct variation rates of position-related parameters and random channel gains in LEO satellite channels, as detailed in Section II, we propose an innovative two-timescale frame structure.
 Downlink positioning operates on the longer position-coherent timescale, while uplink channel estimation and downlink communication occur on the shorter channel gain-coherent timescale.
 Through this framework, each process is scheduled according to its variation rate, maximizing resource utilization while preserving the potential to leverage positioning for efficient channel estimation and communication.
-  • To characterize the LEO satellite-based positioning performance under imperfect inter-satellite clock bias and CFO, we employ misspecified Cramér-Rao bound (MCRB) theory rather than traditional Cramér-Rao bound (CRB) analysis.
-The MCRB framework is specifically designed to derive performance limits for estimation under model mismatch[27].
-Through this theoretical approach, we rigorously analyze how imperfect inter-satellite clock bias and CFO affect time-of-arrival (TOA) and Doppler shift measurements, and systematically derive their impact on positioning accuracy.
-  • We demonstrate the advantage of leveraging the position information of UTs, obtained during the positioning stage, to enhance uplink channel estimation.
-Specifically, we account for the inevitable positioning errors, which introduce model mismatches in the positioning-aided channel estimation (PACE) process.
-By invoking the MCRB theory, we provide a theoretical characterization of the uplink channel estimation performance facilitated by the UTs’ position information.
-  • We propose a distributed beamforming strategy for downlink communication, where each UT is served by the LEO satellites withn the cluster simultaneously.
-In this approach, each LEO satellite independently optimizes its beamformer using only its local CSI with respect to UTs, thereby eliminating the need for a centralized processor and complex inter-satellite coordination mechanisms.
+  • To move beyond idealized synchronization assumptions commonly adopted in prior works, we extend the use of the misspecified Cramér-Rao bound (MCRB) framework to explicitly capture the impact of inter-satellite clock bias and CFO.
+Unlike the classical Cramér-Rao bound (CRB), the MCRB is designed for estimation under model mismatch [38].
+Through this lens, we rigorously quantify how these synchronization-related impairments distort time-of-arrival (TOA) and Doppler measurements and, in turn, degrade positioning accuracy in practical LEO scenarios.
+  • Within the proposed positioning-aided channel estimation (PACE) framework, we show how large-timescale positioning errors propagate into small-timescale channel estimation, creating inherent model mismatch.
+By applying the MCRB, we systematically analyze how uncertainty in the positioning stage translates into performance loss in uplink channel estimation, thereby identifying the robustness margins required for practical position-aided designs.
+  • We propose a cooperative beamforming baseline for downlink communication, where each UT is simultaneously served by all LEO satellites in a cluster.
+We adopt the weighted minimal mean squared error (WMMSE) framework and develop a low-complexity solution that strategically exploits block coordinate descent (BCD), strong duality, and matrix-inversion structures to mitigate the otherwise prohibitive computational burden of multi-satellite cooperation.
+Importantly, this baseline is designed to be computationally practical and to enable a transparent evaluation of the proposed PACE framework by comparing the performance achieved under nominal (PACE-estimated) versus ground-truth channel models.
 This paper is structured as follows.
@@
 Section III examines both downlink positioning and uplink channel estimation performance through MCRB theory, incorporating the PACE.
-A distributed beamforming approach for downlink communications is detailed in Section IV.
-Numerical results are presented in Section V, with conclusions drawn in Section VII.
+A cooperative beamforming approach for downlink communications is detailed in Section IV.
+Numerical results are presented in Section V, with conclusions drawn in Section VI.
 Notations: Scalars are written in regular lowercase, while vectors and matrices are denoted by bold lowercase and bold uppercase letters, respectively.
@@
 1, we consider a scenario where a cluster of $S$ LEO satellites serves $U$ UTs, simultaneously providing downlink positioning and communication services.
+In this paper, we assume that no handovers occur during operation.
+The impact of both intra- and inter-cluster handovers is left for future investigation.
 Each LEO satellite is equipped with a UPA comprising $N=N_{\text{h}}\times N_{\text{v}}$ half-wavelength-spaced antennas, where $N_{\text{h}}$ and $N_{\text{v}}$ denote the numbers of antennas in the horizontal and vertical dimensions, respectively.
 Each UT is equipped with a single antenna.
+[^1] Note that extending single-antenna UTs to multi-antenna setting is a promising direction: receiver spatial resolution can help separate signals with distinct Dopplers and delays and mitigate asynchronous interference[23].
+However, a rigorous treatment would also require modifying the positioning model to incorporate array orientation and angle-related measurements, which is beyond the present scope and will be pursued in future work.
 While accurate CSI is essential for achieving efficient downlink communication, channel estimation is challenging in space-borne communication systems due to the limited link budget and short coherence time.
@@
 Subsequently, we elaborate on a two-timescale property of this model, which can be leveraged to facilitate efficient channel acquisition at LEO satellites, thereby enhancing downlink communications.
+Note that handovers in LEO satellite systems typically occur on the order of tens of seconds to a few minutes, and thus span multiple position-coherent large-timescale frames (see further discussion later).
+Since the proposed PACE framework focuses on the integration of positioning and communication within each frame, we omit explicit handover modeling to avoid introducing an additional control-plane layer and associated redundancy beyond the scope of this work.
+This simplifying assumption is also widely adopted in prior studies such as[24, 19, 15, 16].
 II-A Channel Model
@@
 Consequently, LOS propagation becomes even more dominant.
-Assume the system employs the orthogonal frequency division multiplexing (OFDM) scheme with $K$ subcarriers.
-Let $\Delta f=B/K$ and $T=1/B$ denote the subcarrier spacing and symbol duration, respectively, where $B$ is the signal bandwidth.
-The channel from the $s$-th LEO satellite to the $u$-th UT during the $\ell$-th symbol over the $k$-th subcarrier is expressed as [^2] Note that, though not further indexed for symbolic simplicity, this expression is viewed within one channel-gain coherent interval (as elaborated later) and $\upsilon_{s,u}$, $\tau_{s,u}$, and $\bm{\theta}_{s,u}$ are viewed constant for many such intervals.
+Assume the system employs the OFDM scheme with $K$ subcarriers.
+Let $\Delta f=B/K$ and $T=1/\Delta f$ denote the subcarrier spacing and symbol duration, respectively, where $B$ is the signal bandwidth.
+The channel from the $s$-th LEO satellite to the $u$-th UT during the $\ell$-th OFDM symbol over the $k$-th subcarrier is expressed as [^3] Note that, though not further indexed for symbolic simplicity, this expression is viewed within one channel-gain coherent interval (as elaborated later) and $\upsilon_{s,u}$, $\tau_{s,u}$, and $\bm{\theta}_{s,u}$ are viewed constant for many such intervals.
 $\mathbf{h}_{s,u}\left[\ell,k\right]=\alpha_{s,u}G\left(\theta_{s,u}^{\mathrm{el}}\right)e^{\jmath 2\pi\left(\ell T\upsilon_{s,u}-k\Delta f\tau_{s,u}\right)}\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (1)
-where $\alpha_{s,u}$ is the complex channel gain, $G(\cdot)$ is the antenna radiation gain determined solely by the elevation angle[28], $\tau_{s,u}$ and $\upsilon_{s,u}$ represent the TOA and Doppler shift, respectively, and $\mathbf{a}(\bm{\theta}_{s,u})\in\mathbb{C}^{N}$ is the steering vector at the LEO satellite, with $\bm{\theta}_{s,u}=[\theta_{s,u}^{\mathrm{az}},\theta_{s,u}^{\mathrm{el}}]^{\mathsf{T}}$ being the angle-of-departure (AOD) composed of both azimuth and elevation angles.
-Without loss of generality, we assume the UPA at each satellite is deployed on the XY-plane of its local coordinate system, as illustrated in Fig.
-2.
+where $\alpha_{s,u}$ is the complex channel gain, $G(\cdot)$ is the symmetric antenna radiation pattern determined solely by the elevation angle[39], $\tau_{s,u}$ and $\upsilon_{s,u}$ represent the TOA and Doppler shift, respectively, and $\mathbf{a}(\bm{\theta}_{s,u})\in\mathbb{C}^{N}$ is the steering vector at the LEO satellite, with $\bm{\theta}_{s,u}=[\theta_{s,u}^{\mathrm{az}},\theta_{s,u}^{\mathrm{el}}]^{\mathsf{T}}$ being the angle-of-departure (AOD) composed of both azimuth and elevation angles.
+Without loss of generality, we assume the UPA at each satellite is deployed on the XY-plane of its local coordinate system.
 Hence, the steering vector can be expressed as
@@
 Specifically, the path loss $\beta_{s,u}$, in the dB domain, is characterized as
-[…]
+$20\log_{10}\beta_{s,u}=\beta^{\text{FS}}_{s,u}+\beta^{\text{SF}}_{s,u}+\beta^{\text{CL}}_{s,u}+\beta^{\text{AB}}_{s,u}+\beta^{\text{SC}}_{s,u}\text{ [dB]},$ (4)
 where $\beta^{\text{FS}}_{s,u}$ denotes the free-space path loss, $\beta^{\text{SF}}_{s,u}$ accounts for shadow fading modeled as a Gaussian random variable, $\beta^{\text{CL}}_{s,u}$ represents the clutter loss, $\beta^{\text{AB}}_{s,u}$ captures atmospheric absorption effects, and $\beta^{\text{SC}}_{s,u}$ accounts for attenuation caused by ionospheric or tropospheric scintillation.
-Figure 2: Geometric configuration of the local coordinate systems for LEO satellites, depicting the UPA deployment in the XY-plane.
 II-B Timescale Analysis and Frame Structure
@@
 From (1), we observe that the LEO satellite channel is governed by two categories of parameters:
-  • Large-Timescale Parameters: These parameters are associated with the UT’s position, including the Doppler shift, TOA, and AOD.
-Due to the large geographical scale of satellite communications, they typically change slowly (on timescales of seconds) or can be accurately predicted through tracking mechanisms to reduce estimation overhead[25].
-Although LEO satellites move rapidly, their orbits are typically predictable and well-determined.
-Therefore, while these parameters are related to the fast-varying position of LEO satellites, they can still be considered slow-varying due to their predictability.
+  • Large-Timescale Parameters: These parameters are associated with the positions of the satellites and UTs, and include the Doppler shift, TOA, and AOD.
+Although LEO satellites move rapidly, their orbits are predictable and well-determined.
+Thus, the dominant deterministic Doppler component evolves coherently with other geometry-related parameters such as TOA and AOD.
+As long as the UTs’ positions do not change abruptly (which is typically the case in practice), this predictability remains valid over the position-coherent timescale, which is primarily governed by the update rate of the UTs’ positions rather than the satellites’ trajectories.
+Therefore, it is reasonable to assume that large-timescale parameters vary on the order of seconds or longer in typical scenarios.
   • Small-Timescale Parameters: These parameters consist of random variations in the amplitude and phase of the channel gain.
-In contrast to the slow-varying parameters determined by geometry, these parameters are inherently unpredictable and fluctuate rapidly.
-For instance, tropospheric scintillation can cause phase and amplitude variations in Ka-band LEO satellite channels that fluctuate on sub-second timescales, with the exact rate depending on atmospheric and orbital conditions[31].
-Notably, the Doppler shift caused by the high-speed movement of the satellite induces rapid temporal phase rotation in the channel.
-However, this effect can be largely mitigated through carrier synchronization at the receiver, albeit with minor residual errors [3].
-As a result, its impact on LEO satellite communications is limited.
-Nonetheless, we retain it in the channel model because it can serve as a valuable measurement for positioning, as will be discussed later.
-For positioning purposes, the Doppler shift should be explicitly estimated before being compensated for communications.
+Residual Doppler fluctuations that remain after compensation, due to estimation errors, are also absorbed into this category.
+Unlike the geometry-determined large-timescale parameters, these small-timescale effects are inherently unpredictable and fluctuate rapidly.
+Such variations can arise from tropospheric scintillation, which induces sub-second fluctuations in Ka-band LEO satellite channels depending on atmospheric and orbital conditions [42].
+Moreover, even though the LOS component dominates at Ku/Ka bands, nearby mobile scatterers can still generate multipath components that merge with the LOS path, forming a composite channel gain that exhibits small-scale fading with amplitude and phase variations on millisecond timescales [28, 18].
+Note that a similar dual-timescale effect has been analyzed and exploited in terrestrial systems [43, 27].
+In LEO satellite channels, the Doppler shift induced by the high-speed motion of the satellite causes rapid temporal phase rotations.
+While this effect can be largely mitigated by carrier synchronization at the receiver, small residual errors may remain [8], limiting its impact on communication performance.
+Nevertheless, we retain the Doppler shift in the channel model since it also provides a valuable measurement for positioning, as will be discussed later.
+For positioning, the Doppler shift must be explicitly estimated prior to being compensated for communication purposes.
 II-B2 Two-Timescale Frame Structure
@@
 As shown in Fig.
-3, the duration of a large-timescale frame is determined by the position-coherent time, which indicates the interval after which the UTs’ position information becomes outdated.
-Each large-timescale frame begins with an initial subframe for downlink positioning (where position estimates obtained by UTs are fed back to LEO satellites), followed by multiple small-timescale subframes during which the channel gain remains coherent.
-Each small-timescale subframe is further divided into slots for uplink channel estimation and downlink communication.
-Note that time division duplex (TDD) is adopted to differentiate between uplink and downlink transmissions.
+2, the duration of a large-timescale frame is determined by the position-coherent time, which indicates the interval after which the UTs’ position information becomes outdated.
+Each large-timescale frame begins with an initial subframe for downlink positioning, where UTs’ position estimates can be fed back to LEO satellites [24, 44], followed by multiple small-timescale subframes during which the channel gain remains coherent.
+We would like to emphasize that the feedback carries only low-rate UT position information (a 3D real vector) that remains coherent over the large-timescale interval.
+This leads to negligible overhead and no stringent latency requirements.
+Moreover, the dominant uncertainty is the position estimation error (already modeled in our analysis), under which Fig.
+5 shows that PACE-based channel reconstruction is robust in practical regimes.
+Therefore, explicit feedback-link errors are omitted.
+Note that time division duplex (TDD) [^4] While current NR-NTN bands (e.g., n255/n256 [45]) are mainly frequency division duplex (FDD), our PACE framework relies on uplink–downlink reciprocity and is therefore studied under a TDD assumption.
+Notably, TDD for LEO is explicitly considered by 3rd Generation Partnership Project (3GPP) [41], has operational precedent in systems such as Iridium Next, and is further examined in recent feasibility studies [46].
+Hence, our TDD-based design complements the FDD baseline and provides forward-looking insights into reciprocity-enabled schemes for future NTN systems.
+is adopted to differentiate between uplink and downlink transmissions.
 The proposed approach leverages uplink-downlink channel reciprocity, which applies uplink channel estimation to optimize downlink communication design.
@@
 Alternatively, to achieve equivalent channel estimation accuracy, reduced pilot overhead is required, thus preserving more resources for data transmission.
-In the following, we will first conduct a performance analysis based on the two-timescale PACE and then propose a distributed beamforming method across multiple LEO satellites for downlink communication.
+In the following, we will first conduct a performance analysis based on the two-timescale PACE and then propose a cooperative beamforming method across multiple LEO satellites for downlink communication.
 III Two-timescale Positioning-Aided Channel Estimation
-To characterize the performance limits of the proposed two-timescale PACE, we conduct a two-phase theoretical performance analysis.
+To characterize its performance limits, we conduct a two-phase theoretical performance analysis [^5] For both downlink positioning and uplink channel estimation, we theoretically characterize the performance under model mismatch using the MCRB, with the aim of revealing the potential of the proposed PACE framework in enhancing channel estimation.
+The design of specific estimators is beyond the scope of this paper and is left for future work.
+for PACE.
 In the first phase, we analyze the downlink multi-LEO satellite-based positioning performance, while in the second phase, we assess the uplink channel estimation performance.
-Before analyzing positioning performance, we acknowledge the presence of model mismatches due to hardware imperfections.
-First, low-cost oscillators and inherent hardware defects can introduce phase and frequency errors [32].
-Second, clock biases are inevitable between satellites and UTs as well as between satellites themselves.
-Although a sub-nanosecond clock synchronization can be achieved by deploying GNSS receivers on LEO satellites [33], actual satellite signals remain subject to ionospheric and tropospheric delays during propagation [34].
-These delays, influenced by factors such as temperature, pressure, humidity, and the satellite’s elevation angle relative to the UT, result in unique delays for each satellite-UT pair, effectively introducing an additional clock bias in the signal model.
-Despite these non-idealities, many localization algorithms simplify the model by ignoring such factors to reduce computational complexity.
-For instance, in the scenario considered here, the large beam footprint of LEO satellite communication systems relaxes UT positioning accuracy requirements, while rapid channel variations prioritize model simplicity for timely estimations.
-Accordingly, we adopt a simplified model to estimate unknowns derived from a more complex realistic model, which inherently leads to model mismatches.
-Similarly, in the second phase, namely position information-aided channel estimation, the presence of estimation errors in UT positions also introduces model mismatches.
-Consequently, the performance bound under these model mismatches can be effectively analyzed using the tool of MCRB [27].
-III-A MCRB Basics
-This subsection briefly recaps the fundamentals of MCRB.
-For a parameter $\bar{\bm{\theta}}\in\mathbb{R}^{M\times 1}$, the lower bound (LB) matrix of the mean squared error of a mismatched estimator is provided by[27]
-$\text{LBM}\big{(}\tilde{\bm{\theta}}\big{)}=\underbrace{\mathbf{A}^{-1}_{\tilde{\bm{\theta}}}\mathbf{B}_{\tilde{\bm{\theta}}}\mathbf{A}^{-1}_{\tilde{\bm{\theta}}}}_{\text{MCRB}\left(\tilde{\bm{\theta}}\right)}+\underbrace{\left(\bar{\bm{\theta}}-\tilde{\bm{\theta}}\right)\left(\bar{\bm{\theta}}-\tilde{\bm{\theta}}\right)^{\top}}_{\text{Bias}\left(\tilde{\bm{\theta}}\right)}.$ (5)
-Here, $\tilde{\bm{\theta}}$ denotes the pseudo-true parameter obtained by
-$\tilde{\bm{\theta}}=\arg\underset{\bm{\theta}}{\min\;}\mathcal{D}\left\{\xi_{\text{T}}\left(\mathbf{y};\bar{\bm{\theta}}\right)\left|\right|\xi_{\text{M}}\left(\mathbf{y};\bm{\theta}\right)\right\},$ (6)
-where $\mathcal{D}\{\xi_{\text{T}}\left(\mathbf{y};\bm{\theta}\right)\left|\right|\xi_{\text{M}}(\mathbf{y};\tilde{\bm{\theta}})\}$ denotes the Kullback-Leibler (KL) divergence between the true likelihood function $\xi_{\text{T}}\left(\mathbf{y};\bm{\theta}\right)$ and the mismatched likelihood function $\xi_{\text{M}}(\mathbf{y};\tilde{\bm{\theta}})$.
-In addition, $\mathbf{A}_{\tilde{\bm{\theta}}}$ and $\mathbf{B}_{\tilde{\bm{\theta}}}$ represent two generalized Fisher information matrixs, whose elements in the $i$-th row and the $j$-th column are determined by[27]
-$\left[\mathbf{A}_{\tilde{\bm{\theta}}}\right]_{i,j}=$ $\mathbb{E}_{f_{T}}\left\{\frac{\partial^{2}\ln\xi_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}\partial\left[\bm{\theta}\right]_{j}}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}}\right\}$
-$=$ $2\Re\left\{\left(\frac{\partial^{2}\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}\partial\left[\bm{\theta}\right]_{j}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\bm{\epsilon}\left(\bm{\theta}\right)\right.$ (7)
-$\;\;\;\;\;\;\;\;\;-\left.\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right)\right\}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}},$
-and
-$\left[\mathbf{B}_{\tilde{\bm{\theta}}}\right]_{i,j}=$ $\mathbb{E}_{f_{T}}\left\{\frac{\partial\ln f_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}}\frac{\partial\ln f_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}}\right\}$
-$=$ $4\Re\left\{\bm{\epsilon}\left(\bm{\theta}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right\}\Re\left\{\bm{\epsilon}\left(\bm{\theta}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right\}$
-$+$ $2\Re\left\{\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right)\right\}\bigg{|}_{\bm{\theta}=\tilde{\bm{\theta}}},$ (8)
-respectively.
-Here, $\mathbf{C}_{\text{M}}$ represents the covariance matrix of $\xi_{\text{M}}(\mathbf{y};\bm{\theta})$ and $\bm{\epsilon}(\tilde{\bm{\theta}})=\bm{\kappa}(\bar{\bm{\theta}})-\bm{\mu}(\tilde{\bm{\theta}})$ with $\bm{\kappa}(\bar{\bm{\theta}})$ and $\bm{\mu}(\tilde{\bm{\theta}})$ being the noise-free observations under the true and mismatched models, respectively[27, 35].
-III-B Large-Timescale Downlink Positioning
-III-B1 Signal Model
+Note that model mismatches can arise from various hardware imperfections.
+[^6] While impairments such as phase noise and power amplifier nonlinearities are also relevant at Ku/Ka band[47], we focus on clock bias and CFO because they directly perturb delay and Doppler, introducing deterministic biases across satellites that dominate positioning errors.
+By contrast, phase noise and nonlinearities mainly act as stochastic distortions that reduce effective received power and can, in principle, be mitigated on a single-satellite basis.
+First, clock biases are inevitable between satellites and UTs as well as among satellites themselves.
+Although sub-nanosecond synchronization can be achieved by equipping LEO satellites with GNSS receivers [48], the received signals remain subject to ionospheric and tropospheric delays during propagation [49].
+These delays, shaped by temperature, pressure, humidity, and the satellite’s elevation angle relative to the UT, induce unique offsets for each satellite–UT pair and effectively translate into equivalent clock biases in the signal model.
+Second, low-cost oscillators and inherent hardware impairments can introduce CFO [50].
+Many localization algorithms simplify the underlying model by neglecting certain factors to reduce analytical complexity.
+To capture the impact of this simplification, in the positioning phase we employ a simplified model to estimate unknowns that are actually generated by a more realistic one, which inherently introduces model mismatches.
+Similarly, in the second phase, position information-aided channel estimation, estimation errors in UT positions also lead to mismatches.
+Consequently, the performance bounds under such conditions can be effectively characterized using the MCRB [38, 51].
+III-A Large-Timescale Downlink Positioning
+III-A1 Signal Model
 Let $\mathbf{t}_{s}[\ell,k]\in\mathbb{C}^{U}$ be the pilot (with unit-modulus elements) sent by the $s$-th LEO satellite over the $\ell$-th symbol and the $k$-th subcarrier.
 We assume the satellites are multiplexed in the positioning phase through time-division multiple access (TDMA)[36].
-The received baseband signal at the $u$-th user, from the $s$-th satellite can be expressed as
+The received baseband signal at the $u$-th UT, from the $s$-th satellite can be expressed as
 $y_{s,u}\left[\ell,k\right]=$ $\mathbf{h}_{s,u}^{\mathsf{T}}\left[\ell,k\right]\mathbf{F}_{s}\mathbf{t}_{s}\left[\ell,k\right]+z_{u}\left[\ell,k\right]$
-$=$ $\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}^{\mathsf{T}}\left(\bm{\theta}_{s,u}\right)\mathbf{f}_{s}\left[\ell,k\right]+z_{u}\left[\ell,k\right],$ (9)
+$=$ $\vartheta_{s,u}\gamma\left[\ell,k\right]\mathbf{a}^{\mathsf{T}}\left(\bm{\theta}_{s,u}\right)\mathbf{f}_{s}\left[\ell,k\right]+z_{u}\left[\ell,k\right],$ (5)
 […]
-Let ${\bm{\eta}}_{s,u}=[\acute{\bm{\eta}}_{s,u}^{\mathsf{T}},\grave{\bm{\eta}}_{s,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{6}$ be the unknown channel-domain parameters in signals received from the $s$-th satellite, where $\acute{\bm{\eta}}_{s,u}=[\upsilon_{s,u},\tau_{s,u}]^{\mathsf{T}}\in\mathbb{R}^{2}$ contains the channel-domain parameters we used to localize the user and $\grave{\bm{\eta}}_{s,u}=[\bm{\theta}_{s,u}^{\mathsf{T}},\Re(\alpha_{s,u}),\Im(\alpha_{s,u})]^{\mathsf{T}}\in\mathbb{R}^{4}$ contains the other nuisance parameters that we do not explore.
-Note here the AOD $\bm{\theta}_{s,u}$ is unused for UT positioning due to the limited angular resolution of satellites.
+Let ${\bm{\eta}}_{s,u}=[\acute{\bm{\eta}}_{s,u}^{\mathsf{T}},\grave{\bm{\eta}}_{s,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{6}$ denote the unknown channel-domain parameters in the signal received from the $s$-th satellite, where $\acute{\bm{\eta}}_{s,u}=[\upsilon_{s,u},\tau_{s,u}]^{\mathsf{T}}\in\mathbb{R}^{2}$ are the parameters used for UT positioning, and $\grave{\bm{\eta}}_{s,u}=[\bm{\theta}_{s,u}^{\mathsf{T}},\Re(\vartheta_{s,u}),\Im(\vartheta_{s,u})]^{\mathsf{T}}\in\mathbb{R}^{4}$ are treated as nuisance parameters.
+Delay $\tau_{s,u}$ and Doppler $\upsilon_{s,u}$ are directly determined by the UT’s position and velocity relative to the satellites, making them the most reliable observables for positioning.
+In contrast, the AOD $\bm{\theta}_{s,u}$ is not exploited here due to the limited angular resolution of LEO satellites operating at high altitudes, which makes AOD estimates unreliable unless extremely large arrays are employed.
+Similarly, the channel gain $\vartheta_{s,u}$ is heavily influenced by environmental variability and hardware impairments and thus provides little stable geometric information for positioning, except in fingerprinting-based approaches [40].
 We further concatenate all the utilized channel-domain parameters collected from $S$ satellites as $\acute{\bm{\eta}}_{u}=[\acute{\bm{\eta}}_{1,u}^{\mathsf{T}},\ldots,\acute{\bm{\eta}}_{S,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{2S}$.
 Therefore, in the large-timescale downlink positioning, we first estimate $\acute{\bm{\eta}}_{u}$ based on $y_{s,u}[\ell,k]$, $s=1,\dots,S$, $\ell=1,\dots,L_{\text{p}}$, $k=1,\dots,K$, where $L_{\text{p}}$ denotes the number of pilot symbols for downlink positioning.
-Afterwards, we estimate the $u$-th user’s position, clock biases, and CFOs.
-III-B2 The FIM of ${\bm{\eta}}_{s,u}$
-Let ${\mu}_{s,u}[\ell,k]$ denote the noise-free version of $y_{s,u}[\ell,k]$ and $\mathbf{G}_{s,u}[\ell,k]=[\mathbf{g}_{v}[\ell,k],\mathbf{g}_{\tau}[\ell,k],\mathbf{g}_{\theta_{\mathrm{az}}}[\ell,k],\mathbf{g}_{\theta_{\mathrm{el}}}[\ell,k],\mathbf{g}_{\alpha,\mathrm{R}}[\ell,k],\mathbf{g}_{\alpha,\mathrm{I}}[\ell,k]]\in\mathbb{C}^{N\times 6}$.
+Afterwards, we estimate the $u$-th UT’s position, clock biases, and CFOs.
+III-A2 The FIM of ${\bm{\eta}}_{s,u}$
+Let ${\mu}_{s,u}[\ell,k]$ denote the noise-free version of $y_{s,u}[\ell,k]$ and $\mathbf{G}_{s,u}[\ell,k]=[\mathbf{g}_{v}[\ell,k],\mathbf{g}_{\tau}[\ell,k],\mathbf{g}_{\theta_{\mathrm{az}}}[\ell,k],\mathbf{g}_{\theta_{\mathrm{el}}}[\ell,k],\mathbf{g}_{\vartheta,\mathrm{R}}[\ell,k],\mathbf{g}_{\vartheta,\mathrm{I}}[\ell,k]]\in\mathbb{C}^{N\times 6}$.
 Here, the columns fulfill
-$\mathbf{g}_{v}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\left(\jmath 2\pi\ell T\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10a)
-$\mathbf{g}_{\tau}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\left(-\jmath 2\pi k\Delta f\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10b)
-$\mathbf{g}_{\theta_{\mathrm{az}}}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{\mathrm{az}}},$ (10c)
-$\mathbf{g}_{\theta_{\mathrm{el}}}\left[\ell,k\right]$ $=\alpha_{s,u}\left(\frac{\partial\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]}{\partial\theta_{s,u}^{\mathrm{el}}}\mathbf{a}\left(\bm{\theta}_{s,u}\right)\right.$ (10d)
-$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left.\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{s,u}^{\mathrm{el}}}\right),$
-$\mathbf{g}_{\alpha,\mathrm{R}}\left[\ell,k\right]$ $=\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10e)
-$\mathbf{g}_{\alpha,\mathrm{I}}\left[\ell,k\right]$ $=\jmath\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right).$ (10f)
-Based on the signal received from the $s$-th satellite, the FIM of ${\bm{\eta}}_{s,u}$ can be computed by the Slepian-Bangs formula as
+$\mathbf{g}_{v}\left[\ell,k\right]$ $=\vartheta_{s,u}\gamma\left[\ell,k\right]\left(\jmath 2\pi\ell T\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (6a)
+$\mathbf{g}_{\tau}\left[\ell,k\right]$ $=\vartheta_{s,u}\gamma\left[\ell,k\right]\left(-\jmath 2\pi k\Delta f\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (6b)
+$\mathbf{g}_{\theta_{\mathrm{az}}}\left[\ell,k\right]$ $=\vartheta_{s,u}\gamma\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{s,u}^{\mathrm{az}}},$ (6c)
+$\mathbf{g}_{\theta_{\mathrm{el}}}\left[\ell,k\right]$ $=\vartheta_{s,u}\gamma\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{s,u}^{\mathrm{el}}},$ (6d)
+$\mathbf{g}_{\vartheta,\mathrm{R}}\left[\ell,k\right]$ $=\gamma\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (6e)
+$\mathbf{g}_{\vartheta,\mathrm{I}}\left[\ell,k\right]$ $=\jmath\gamma\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right).$ (6f)
+Based on the signal received from the $s$-th satellite, the Fisher information matrix (FIM) of ${\bm{\eta}}_{s,u}$ can be computed by the Slepian-Bangs formula as
 $\mathbf{J}_{s}({\bm{\eta}}_{s,u})$
@@
 $=$ $\frac{2}{N_{0}\Delta f}\sum_{\ell=1}^{L_{\text{p}}}\sum_{k=1}^{K}\Re\big\{\mathbf{G}_{s,u}^{\mathsf{T}}\left[\ell,k\right]\mathbf{f}_{s}\left[\ell,k\right]\mathbf{f}_{s}\left[\ell,k\right]^{\mathsf{H}}\mathbf{G}_{s,u}^{*}\left[\ell,k\right]\big\}.$
-III-B3 The FIM of $\acute{\bm{\eta}}_{u}$
+III-A3 The FIM of $\acute{\bm{\eta}}_{u}$
 Partitioning $\mathbf{J}_{s}(\bm{\eta}_{s,u})=[\mathbf{X},\mathbf{Y};\mathbf{Y}^{\mathsf{T}},\mathbf{Z}]$, where $\mathbf{X}\in\mathbb{R}^{2\times 2}$, we can compute the FIM of the total utilized channel-domain parameters $\acute{\bm{\eta}}_{s,u}$ as $\mathbf{J}_{s}\left(\acute{\bm{\eta}}_{s,u}\right)=\mathbf{X}-\mathbf{Y}\mathbf{Z}^{-1}\mathbf{Y}^{\mathsf{T}}$.
@@
 $\mathbf{J}\left(\acute{\bm{\eta}}_{u}\right)=\text{blkdiag}\big\{\mathbf{J}_{1}\left(\acute{\bm{\eta}}_{1,u}\right),\ldots,\mathbf{J}_{S}\left(\acute{\bm{\eta}}_{S,u}\right)\big\}.$ (8)
-Based on this, we can infer that the achievable lowest estimation error on $\acute{\bm{\eta}}_{u}$ regardless of the adopted algorithms is characterized by the Gaussian distribution $\mathcal{N}\big{(}\mathbf{0},\mathbf{J}_{s}^{-1}(\acute{\bm{\eta}}_{u})\big{)}$.
+Based on this, we can infer that the achievable lowest estimation error on $\acute{\bm{\eta}}_{u}$ regardless of the adopted algorithms is characterized by the Gaussian distribution $\mathcal{N}\big(\mathbf{0},\mathbf{J}^{-1}(\acute{\bm{\eta}}_{u})\big)$.
 Any algorithm achieving this error covariance matrix is called an efficient estimator.
-III-B4 MCRB and LB Derivation
+III-A4 MCRB and LB Derivation
 Suppose an efficient estimator is utilized to estimate $\acute{\bm{\eta}}_{u}$.
@@
 Here, as the velocity of LEO satellite is usually much higher than that of the UT, we ignore the velocity contribution from the UT side.
-Based on the underlying geometric relationship, the forward model between the entries of $\acute{\bm{\eta}}_{u}$ and unknown position of the $u$-th user can be expressed as
-$\upsilon_{s,u}$ $=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u}+\Delta\delta_{s,u},$ (13a)
-$\tau_{s,u}$ $=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}+\Delta b_{s,u},$ (13b)
+This assumption is justified since LEO satellites move at around 7.6 km/s, whereas UTs typically move at only a few m/s (pedestrians) to a few hundred m/s (vehicles/aircraft), making their velocity contribution two to three orders of magnitude smaller and thus negligible for the Doppler term.
+Based on the underlying geometric relationship, the forward model between the entries of $\acute{\bm{\eta}}_{u}$ and unknown position of the $u$-th UT can be expressed as
+$\upsilon_{s,u}=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u}+\Delta\delta_{s,u}$ (9)
+and
+$\tau_{s,u}=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}+\Delta b_{s,u},$ (10)
 which we refer to as the true model.
+Here, $\mathbf{p}_{u}\in\mathbb{R}^{3}$ and $\mathbf{p}_{s}\in\mathbb{R}^{3}$ represent the positions of the $u$-th UT and the $s$-th satellite [^7] Note that LEO satellites’ positions are assumed precisely known at the UTs.
+This information can be obtained from the ephemerides contained in the two-line element files, which are updated daily by the North American Aerospace Defense Command (NORAD) [36]., respectively.
 In real applications, we often ignore these clock bias and CFO differences, i.e., assuming all satellite-UT pairs have the same clock bias and CFO, for algorithmic simplicity.
-To account for the impact of such model mismatch, here we derive the MCRB.
+To account for the impact of such model mismatch, we derive the MCRB, with its fundamentals detailed in Appendix A for completeness.
 When neglecting the difference on each clock bias and CFO, the mismatched model is given by
-$\upsilon_{s,u}$ $=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u},$ (14a)
-$\tau_{s,u}$ $=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}.$ (14b)
-Now, we suppose a mismatched estimator that aims to estimate the unknown position of the $u$-th user based on the realistic observation $\acute{\bm{\eta}}_{u}$ and simplified model (14) .
-the total unknown parameters in the mismatch model can be concatenated as $\mathbf{r}_{u}=[\mathbf{p}_{u}^{\mathsf{T}},\delta_{u},b_{u}]^{\mathsf{T}}\in\mathbb{R}^{5}$.
-Based on the true model (13) and mismatched model (14), we can respectively express the true likelihood function $\xi_{\mathrm{T}}$ and the mismatched likelihood function $\xi_{\mathrm{M}}$ as
-$\ln{\xi_{\mathrm{T}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto\big{(}\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)}^{\mathsf{T}}\bm{\Sigma}^{-1}\big{(}\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)},$ (15)
+$\upsilon_{s,u}=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u}$ (11)
 and
-$\ln{\xi_{\mathrm{M}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto\big{(}\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)}^{\mathsf{T}}\bm{\Sigma}^{-1}\big{(}\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)},$ (16)
-respectively, where the function $\bar{\mathbf{f}}(\cdot)$ maps $\mathbf{r}_{u}$ to $\acute{\bm{\eta}}_{u}$ according to (13), and the function $\tilde{\mathbf{f}}(\cdot)$ maps $\mathbf{r}_{u}$ to $\acute{\bm{\eta}}_{u}$ according to (14).
-As mentioned in Section III-B3, $\acute{\bm{\eta}}_{u}$ is the estimated channel-domain parameters through an efficient estimator, i.e., $\acute{\bm{\eta}}_{u}\sim\mathcal{N}(\bar{\mathbf{f}}(\mathbf{r}_{u}),\bm{\Sigma})$, where $\bm{\Sigma}=\mathbf{J}^{-1}(\acute{\bm{\eta}}_{u})$.
+$\tau_{s,u}=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}.$ (12)
+Now, we suppose an estimator that aims to estimate the unknown position of the $u$-th UT based on the realistic observation $\acute{\bm{\eta}}_{u}$ and mismatched model, the total unknown parameters in the mismatch model can be concatenated as $\mathbf{r}_{u}=[\mathbf{p}_{u}^{\mathsf{T}},\delta_{u},b_{u}]^{\mathsf{T}}\in\mathbb{R}^{5}$.
+Based on the true model and mismatched model, we can respectively express the true likelihood function $\xi_{\mathrm{T}}$ and the mismatched likelihood function $\xi_{\mathrm{M}}$ as
+${\xi_{\mathrm{T}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto e^{\big(\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big)^{\mathsf{T}}\bm{\Sigma}^{-1}\big(\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big)}$ (13)
+and
+${\xi_{\mathrm{M}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto e^{\big(\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big)^{\mathsf{T}}\bm{\Sigma}^{-1}\big(\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big)},$ (14)
+[…]
   • Pseudo-True Parameter: According to (40), the pseudo-true parameter, denoted by $\tilde{\mathbf{r}}_{u}$, is given by
@@
 The above problem can be solved using gradient descent with backtracking line search, wherein the true parameters serve as the initial point[52].
-  • MCRB: According to (III-A) and (III-A), the generalized FIMs, i.e., $\mathbf{A}_{\tilde{\mathbf{r}}_{u}}$ and $\mathbf{B}_{\tilde{\mathbf{r}}_{u}}$, can be computed by determining the first- and second-order derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to the elements of $\mathbf{r}_{u}$, as detailed in Appendix A.
+  • MCRB: According to Appendix A, the generalized FIMs, i.e., $\mathbf{A}_{\tilde{\mathbf{r}}_{u}}$ and $\mathbf{B}_{\tilde{\mathbf{r}}_{u}}$, can be computed by determining the first- and second-order derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to the elements of $\mathbf{r}_{u}$, as detailed in Appendix B.
 Subsequently, the MCRB is expressed as
     $\text{MCRB}\left(\tilde{\mathbf{r}}_{u}\right)=\mathbf{A}_{\tilde{\mathbf{r}}_{u}}^{-1}\mathbf{B}_{\tilde{\mathbf{r}}_{u}}\mathbf{A}_{\tilde{\mathbf{r}}_{u}}^{-1}.$ (16)
-  • LB: Finally, the estimation error LB matrix can be computed according to (5) as
+  • LB: Finally, the estimation error lower bound (LB) matrix can be computed as
     $\text{LBM}\left(\tilde{\mathbf{r}}_{u}\right)=\text{MCRB}\left(\tilde{\mathbf{r}}_{u}\right)+\text{Bias}\left(\tilde{\mathbf{r}}_{u}\right),$ (17)
@@
     $\text{LB}\left(\mathbf{p}_{u}\right)=\sqrt{\text{Tr}\big(\left[\text{LBM}\left(\tilde{\mathbf{r}}_{u}\right)\right]_{1:3,1:3}\big)}.$ (18)
-III-C Small-Timescale Uplink Channel Estimation
-III-C1 Signal Model
+III-B Small-Timescale Uplink Channel Estimation
+III-B1 Signal Model
 After obtaining the positions of UTs, these large-timescale parameters are fed back to the LEO satellites.
@@
 $\mathbf{y}_{s,u}\left[\ell,k\right]=\underbrace{\vartheta_{s,u}e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\mathbf{d}_{s,u}[\ell,k]}_{\bar{\mathbf{g}}_{s,u}\left[\ell,k\right]}+\mathbf{z}_{s}\left[\ell,k\right],$ (19)
-where $\vartheta_{s,u}=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})$ and $\mathbf{z}_{s}[\ell,k]\sim\mathcal{CN}(\mathbf{0},N_{0}\Delta f\mathbf{I}_{N})$ represents the AWGN.
+where $\mathbf{z}_{s}[\ell,k]\sim\mathcal{CN}(\mathbf{0},N_{0}\Delta f\mathbf{I}_{N})$ represents the AWGN.
 We observe that $\mathbf{d}_{s,u}[\ell,k]$ is directly determined by the UTs’ position feedback.
 Consequently, instead of estimating the full channel vector, it suffices to estimate only two scalars, i.e., the channel gain $\vartheta_{s,u}$ and the TOA $\tau_{s,u}$.
+[^8] Note that although the TOA $\tau_{s,u}$ has already been estimated during the downlink positioning process, it will nevertheless be estimated during uplink channel estimation as part of standard OFDM processing [53].
+This is necessary unless the UT and the LEO satellite are perfectly time-synchronized and the propagation delay remains precisely known, which are hardly satisfied in the highly dynamic LEO satellite environment.
 This significantly simplifies the uplink channel estimation task.
@@
 Here, $\tilde{\bm{\theta}}_{s,u}$ represents the calculated angle-of-arrival (AOA) based on the estimated $\tilde{\mathbf{p}}_{u}$ and the geometrical relationship, given by
-[…]
+$\tilde{\theta}_{s,u}^{\mathrm{az}}$ $=\text{atan2}\left(\left[\tilde{\mathbf{p}}_{u}-\mathbf{p}_{s}\right]_{2},\left[\tilde{\mathbf{p}}_{u}-\mathbf{p}_{s}\right]_{1}\right),$ (21a)
 $\tilde{\theta}_{s,u}^{\mathrm{el}}$ $=\text{asin}\Big(\frac{\left[\left(\tilde{\mathbf{p}}_{u}-\mathbf{p}_{s}\right)\right]_{3}}{\left\|\tilde{\mathbf{p}}_{u}-\mathbf{p}_{s}\right\|}\Big).$ (21b)
-III-C2 MCRB and LB Derivation
+III-B2 MCRB and LB Derivation
 We denote the total temporal and spectral observation as $\mathbf{y}_{s,u}=[\mathbf{y}_{s,u}^{\mathsf{T}}[1,1],\ldots,\mathbf{y}_{s,u}^{\mathsf{T}}[1,K],\ldots,\mathbf{y}_{s,u}^{\mathsf{T}}[L_{\text{c}},K]]^{\mathsf{T}}\in\mathbb{C}^{L_{\text{c}}KN}$, where $L_{\text{c}}$ represents the number of pilot symbols used for uplink channel estimation.
@@
 The likelihood functions under the true and mismatched models are given by
-$\ln{\xi_{\mathrm{T}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto\left\|\mathbf{y}_{s,u}-\bar{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (24)
+${\xi_{\mathrm{T}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto e^{\left\|\mathbf{y}_{s,u}-\bar{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2}}$ (22)
 and
-$\ln{\xi_{\mathrm{M}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto\left\|\mathbf{y}_{s,u}-\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (25)
+${\xi_{\mathrm{M}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto e^{\left\|\mathbf{y}_{s,u}-\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2}},$ (23)
 respectively, where we define $\bar{\mathbf{g}}(\bm{\eta}_{s,u})=[\bar{\mathbf{g}}_{s,u}^{\mathsf{T}}[1,1],\ldots,\bar{\mathbf{g}}_{s,u}^{\mathsf{T}}[1,K],\ldots,\bar{\mathbf{g}}_{s,u}^{\mathsf{T}}[L_{\text{c}},K]]^{\mathsf{T}}\in\mathbb{C}^{L_{\text{c}}KN}$ and $\tilde{\mathbf{g}}(\bm{\eta}_{s,u})=[\tilde{\mathbf{g}}_{s,u}^{\mathsf{T}}[1,1],\ldots,\tilde{\mathbf{g}}_{s,u}^{\mathsf{T}}[1,K],\ldots,\tilde{\mathbf{g}}_{s,u}^{\mathsf{T}}[L_{\text{c}},K]]^{\mathsf{T}}\in\mathbb{C}^{L_{\text{c}}KN}$.
-  • Pseudo-True Parameter: According to (6), the pseudo-true parameter, denoted by $\tilde{\bm{\eta}}_{s,u}$, is derived as
+  • Pseudo-True Parameter: According to Appendix A, the pseudo-true parameter, denoted by $\tilde{\bm{\eta}}_{s,u}$, is derived as
     $\tilde{\bm{\eta}}_{s,u}$ $=\arg\min_{\bm{\eta}_{s,u}}\ \mathcal{D}\big\{{\xi_{\mathrm{T}}}\left(\mathbf{y}_{s,u};\bar{\bm{\eta}}_{s,u}\right)\|{\xi_{\mathrm{M}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\big\},$
-    $=\arg\min_{\bm{\eta}_{s,u}}\ \left\|\bar{\mathbf{q}}\left(\bar{\bm{\eta}}_{s,u}\right)-\tilde{\mathbf{q}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (26)
+    $=\arg\min_{\bm{\eta}_{s,u}}\ \left\|\bar{\mathbf{g}}\left(\bar{\bm{\eta}}_{s,u}\right)-\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (24)
     where the bar above the parameter emphasizes the ground-truth value.
 Once again, by employing gradient descent and treating truth parameter as initial point, the above problem can be efficiently solved [52].
-  • MCRB: The generalized FIMs, i.e., $\mathbf{A}_{\tilde{\bm{\eta}}_{s,u}}$ and $\mathbf{B}_{\tilde{\bm{\eta}}_{s,u}}$, are obtained by calculating the first- and second-order derivatives of $\tilde{\mathbf{q}}\left(\bm{\eta}_{s,u}\right)$ with respect to the elements of $\bm{\eta}_{s,u}$, as described in Appendix B.
+  • MCRB: The generalized FIMs, i.e., $\mathbf{A}_{\tilde{\bm{\eta}}_{s,u}}$ and $\mathbf{B}_{\tilde{\bm{\eta}}_{s,u}}$, are obtained by calculating the first- and second-order derivatives of $\tilde{\mathbf{q}}\left(\bm{\eta}_{s,u}\right)$ with respect to the elements of $\bm{\eta}_{s,u}$, as described in Appendix C.
 Based on these, the MCRB is given by
@@
 For an efficient estimator, the LB for the expected RMSE of the channel gain $\vartheta_{s,u}$ and the TOA $\tau_{s,u}$, in the presence of model mismatch, are expressed as
-    $\text{LB}\left(\vartheta_{s,u}\right)=\sqrt{\text{Tr}\big{(}\left[\text{LBM}\left(\tilde{\bm{\eta}}_{s,u}\right)\right]_{1:2,1:2}\big{)}},$ (29)
+    $\text{LB}\left(\vartheta_{s,u}\right)=\sqrt{\text{Tr}\big(\left[\text{LBM}\left(\tilde{\bm{\eta}}_{s,u}\right)\right]_{1:2,1:2}\big)}$ (27)
     and
@@
     respectively.
-IV Distributed Beamforming for Downlink Communications
+IV Cooperative Beamforming for Downlink Communications
 The proposed two-timescale PACE enables channel reconstruction between each LEO satellite and all UTs, enabling CSI-based beamforming for downlink communication.
-To address the limited link budget in single-satellite downlink scenarios, we introduce distributed multi-LEO satellite beamforming where each UT is served by the same satellite cluster.
-This distributed approach eliminates the need for inter-satellite CSI sharing, as beamforming decisions are made locally at each satellite using only its CSI, rather than requiring consolidated global CSI, which significantly reduces system complexity and latency.
-IV-A Signal Model
-Without centralized coordination, we manage inter-satellite interference by allocating orthogonal frequency bands to different LEO satellites.
-The total subcarrier set $\Phi=\{1,2,\ldots,K\}$ is partitioned into $S$ equal subsets $\Phi_{1},\Phi_{2},\ldots,\Phi_{S}$ corresponding to $S$ LEO satellites.
-For each satellite $s$, its subset is defined as $\Phi_{s}=\{(s-1)\bar{K}+1,(s-1)\bar{K}+1,\ldots,s\bar{K}\}$, where $\bar{K}=K/S$.
-For the $s$-th LEO satellite transmitting over the $k$-th subcarrier ($k\in\Phi_{s}$), the signal is given by
+To address the limited link budget in single-satellite downlink scenarios, we introduce multi-LEO satellite cooperative beamforming, where each UT is served by the satellites within the cluster simultaneously.
+Note that the cooperative beamforming optimization developed in this section is based on a nominal signal model, which treats the channel parameters estimated via PACE as true and assumes perfect delay/Doppler compensation across satellites.
+In contrast, the communication rate in the simulations is evaluated using the ground-truth channel parameters under imperfect compensation.
+This separation enables a direct assessment of the effectiveness of PACE by examining the performance achieved by beamformers optimized with nominal (PACE-estimated) parameters.
+The smaller the performance gap between nominal- and ground-truth-optimized beamformers, the more reliable PACE can be considered.
+Developing robust beamforming strategies that explicitly account for CSI uncertainty (e.g., [27, 54, 55]) and/or asynchronous interference due to imperfect compensation (e.g., [23]) represents an important but orthogonal research direction, which is left for future work.
+IV-A Nominal Signal Model
+In the nominal signal model, we assume perfect delay and Doppler compensation across all satellites.
+Under this assumption, the nominal channel, denoted by $\tilde{\mathbf{h}}_{s,u}=\tilde{\vartheta}_{s,u}\mathbf{a}(\tilde{\bm{\theta}}_{s,u})$, is frequency-flat, where the tilde denotes estimated values.
+This allows subcarrier-coherent beamforming, where the same beamformer is applied across all subcarriers, thereby reducing optimization complexity.
+Specifically, for the $s$-th LEO satellite transmitting over the $k$-th subcarrier, the signal is
 $\mathbf{x}_{s}\left[k\right]=\mathbf{W}_{s}\mathbf{s}\left[k\right],$ (29)
-where $\mathbf{W}_{s}=[\mathbf{w}_{s,1},\mathbf{w}_{s,2},\ldots,\mathbf{w}_{s,U}]\in\mathbb{C}^{N\times U}$ represents the beamformer [^3] We adopt subcarrier-coherent beamforming, where the beamformer is consistent across subcarriers, with only amplitude adjustments allowed.
-This approach strikes a balance between performance and computational efficiency, which is critical given the limited onboard processing capabilities of LEO satellites.
-Although less flexible than subcarrier-specific beamforming, it substantially reduces optimization complexity in systems with a large number of subcarriers.
-The framework can be extended to subcarrier-specific beamforming, albeit at the expense of increased computational complexity.
-for the $s$-th LEO satellite, and $\mathbf{s}\left[k\right]\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{U})$ denotes the data streams transmitted to $U$ UTs over the $k$-th subcarrier.
-The received signal at the $u$-th UT from the $s$-th LEO satellite over the $k$-th subcarrier can be represented as
-$y_{u}\left[k\right]=$ $\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,u}s_{u}\left[k\right]$
-$+\underbrace{\sum_{m=1,m\neq u}^{U}\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,m}s_{m}\left[k\right]}_{\text{Inter-user interference}}+z_{u}\left[k\right],$ (32)
-where $\mathbf{h}_{s,u}[k]=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\mathbf{a}(\bm{\theta}_{s,u})$ denotes the channel vector incorporating Doppler shift compensation.
-The term $z_{u}[k]\sim\mathcal{CN}(\mathbf{0},N_{0}\Delta f)$ represents the AWGN.
-The achievable communication rate for the $u$-th UT, aggregated across all satellites and assigned subcarriers, is given by
-$R_{u}=\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}\left[k\right],$ (33)
-where $R_{s,u}\left[k\right]$ represents the communication rate between the $s$-th LEO satellite and the $u$-th UT over the $k$-th subcarrier, calculated as
-$R_{s,u}\left[k\right]=\log_{2}\left(1+\frac{\left|\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,u}\right|^{2}}{I_{s,u}\left[k\right]+N_{0}\Delta f}\right),$ (34)
-where $I_{s,u}[k]=\sum_{m=1,m\neq u}^{U}|\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,m}|^{2}$.
+where $\mathbf{W}_{s}=[\mathbf{w}_{s,1},\mathbf{w}_{s,2},\ldots,\mathbf{w}_{s,U}]\in\mathbb{C}^{N\times U}$ represents the beamformer for the $s$-th LEO satellite, and $\mathbf{s}\left[k\right]\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{U})$ denotes the data streams transmitted to the $U$ UTs over the $k$-th subcarrier.
+The received signal at the $u$-th UT over the $k$-th subcarrier is then
+$y_{u}\left[k\right]=\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}s_{u}\left[k\right]+\underbrace{\sum_{m\neq u}^{U}\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}s_{m}\left[k\right]}_{\text{Inter-user interference}}+z_{u}\left[k\right],$ (30)
+where $z_{u}[k]\sim\mathcal{CN}(0,N_{0}\Delta f)$ denotes the AWGN.
+The nominal achievable communication rate for the $u$-th UT, aggregated across all subcarriers, is given by
+$\tilde{R}_{u}=B\log_{2}\left(1+\frac{\left|\sum\limits_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}\right|^{2}}{\sum\limits_{m\neq u}^{U}\left|\sum\limits_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}\right|^{2}+N_{0}\Delta f}\right).$ (31)
 IV-B Problem Formulation
 We aim to optimize the beamformers at each LEO satellite to maximize the sum rate of all UTs. The optimization problem is formulated as follows
-$\mathop{\max}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}R_{u}=\sum_{u=1}^{U}\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}[k]$
-${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}},\;\forall s,$ (35)
+$\mathop{\max}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\tilde{R}_{u}$
+${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{K},\;\forall s,$ (32)
 where $P_{s}$ denotes the power budget at the $s$-th LEO satellite.
-Note that, because $\sum_{u=1}^{U}\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}[k]=\sum_{s=1}^{S}\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}R_{s,u}[k]$, the problem in (IV-B) is inherently distributed.
-Consequently, a subproblem corresponding to the $s$-th LEO satellite can be expressed as
-$\mathop{\max}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}R_{s,u}[k]$
-${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}}.$ (36)
-By solving (IV-B) for each LEO satellite, the original problem in (IV-B) can be effectively addressed.
-IV-C Proposed Distributed Beamforming Approach
-In the following, we employ the weighted sum mean squared error minimization (WMMSE) optimization framework [36] to solve (IV-B) iteratively.
-Specifically, by introducing the auxiliary variables $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$, where $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$ serve as the MMSE equalizer and weight, respectively [36], (IV-B) can be equivalently reformulated as
-$\mathop{\max}\limits_{\mu_{s,u}[k],\omega_{s,u}[k],\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}\log_{2}\omega_{s,u}[k]-\omega_{s,u}[k]\Psi_{s,u}[k]$
-${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}},$ (37)
-where $\Psi_{s,u}[k]=|1-\mu_{s,u}[k]\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,u}|^{2}+|\mu_{s,u}[k]|^{2}(I_{s,u}[k]+N_{0}\Delta f)$.
-The optimization problem in (IV-C) is solved by iteratively updating $\mathbf{W}_{s}$, $\omega_{s,u}[k]$, and $\mu_{s,u}[k]$, as detailed below.
-IV-C1 Update $\mu_{s,u}[k]$
-With $\omega_{s,u}[k]$ and $\mathbf{W}_{s}$ fixed, the optimal $\mu_{s,u}[k]$ is derived by setting $\partial\Psi_{s,u}[k]/\partial\mu_{s,u}[k]=0$, resulting in
-$\mu_{s,u}[k]=\frac{\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,u}}{I_{s,u}[k]+N_{0}\Delta f}.$ (38)
-IV-C2 Update $\omega_{s,u}[k]$
-With $\mu_{s,u}[k]$ and $\mathbf{W}_{s}$ fixed, the optimal $\omega_{s,u}[k]$ is obtained by setting $\partial\Psi_{s,u}[k]/\partial\omega_{s,u}[k]=0$, leading to
-$\omega_{s,u}[k]=\frac{1}{\Psi_{s,u}[k]}.$ (39)
+Algorithm 1 Proposed Algorithm for Solving (IV-B)
+1:Initialize: $\mathbf{W}_{s},\forall s$;
+2:repeat
+3:  Update $\mu_{u}$ using (34);
+4:  Update $\omega_{u}$ using (35);
+5:  for $s=1:S$ do
+6:   Obtain eigenvalue decomposition of $\mathbf{R}_{s}$;
+7:   Determine $\lambda$ by Golden-section search over (53);
+8:   Determine $\mathbf{w}_{s}$ via (54) and reshape to obtain $\mathbf{W}_{s}$;
+9:  end for
+10:until the reduction ratio of the objective value falls below a predefined threshold or a predetermined iteration number is reached;
+11:Output: $\mathbf{W}_{s},\forall s$.
+IV-C Proposed Cooperative Beamforming Approach
+In the following, we employ the WMMSE optimization framework [56] to solve (IV-B) iteratively.
+Specifically, by introducing the auxiliary variables $\mu_{u}$ and $\omega_{u}$, where $\mu_{u}$ and $\omega_{u}$ serve as the MMSE equalizer and weight, respectively [56], (IV-B) can be equivalently reformulated as
+$\mathop{\max}\limits_{\mu_{u},\omega_{u},\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\left(\ln\omega_{u}-\omega_{u}\Psi_{u}\right)$
+${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{K},\;\forall s,$ (33)
+where $\Psi_{u}=|1-\mu_{u}\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}|^{2}+|\mu_{u}|^{2}(\sum_{m\neq u}^{U}|\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}|^{2}+N_{0}\Delta f)$.
+The optimization problem in (IV-C) is solved by iteratively updating $\mathbf{W}_{s}$, $\omega_{u}$, and $\mu_{u}$, as detailed below.
+IV-C1 Update $\mu_{u}$
+With $\omega_{u}$ and $\mathbf{W}_{s}$ fixed, the optimal $\mu_{u}$ is derived by setting $\partial\Psi_{u}/\partial\mu_{u}=0$, resulting in
+$\mu_{u}=\frac{\left(\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}\right)^{*}}{\sum_{m=1}^{U}\left|\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}\right|^{2}+N_{0}\Delta f}.$ (34)
+IV-C2 Update $\omega_{u}$
+With $\mu_{u}$ and $\mathbf{W}_{s}$ fixed, the optimal $\omega_{u}$ is obtained by setting $\partial\Psi_{u}/\partial\omega_{u}=0$, leading to
+$\omega_{u}=\frac{1}{\Psi_{u}}.$ (35)
 IV-C3 Update $\mathbf{W}_{s}$
-With $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$ fixed, the optimization problem to find the optimal $\mathbf{W}_{s}$ is formulated as
-$\mathop{\min}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}\Psi_{s,u}[k]$
-${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}}.$ (40)
-Here, $\Psi_{s,u}[k]$ can be expressed as the sum of a positive semi-definite quadratic form and a linear term.
-Consequently, (IV-C3) is a convex quadratically constrained quadratic program (QCQP) problem with respect to $\mathbf{W}_{s}$, which can be efficiently solved via convex optimization toolboxes such as CVX.
-The process of solving (IV-B) is outlined in Algorithm 1.
-The initial beamformer $\mathbf{W}_{s}$ for the $s$-th LEO satellite can be constructed using maximum ratio transmission (MRT), which aligns the beamformer with the reconstructed channels (by excluding the frequency-dependent delay-induced phase rotation).
-Specifically, we have $\mathbf{W}_{s}=\zeta_{s}[\hat{\vartheta}_{s,1}^{*}\mathbf{a}^{*}(\hat{\bm{\theta}}_{s,1}),\ldots,\hat{\vartheta}_{s,U}^{*}\mathbf{a}^{*}(\hat{\bm{\theta}}_{s,U})]$, where the hat notation indicates estimated values and $\zeta_{s}$ is a normalization factor that ensures the fulfillment of the power budget.
-Since the WMMSE optimization framework is utilized to solve each subproblem, i.e., (IV-B), its convergence is well-established [36].
-The computational complexity of each iteration is primarily determined by solving (IV-C3), which, due to its QCQP nature, is characterized by $\mathcal{O}(N^{3}U^{3})$.
-Algorithm 1 Algorithm for Solving (IV-B)
-1:Initialize: $\mathbf{W}_{s},\forall s$;
-2:for $s=1:S$ do
-3:     repeat
-4:         Update $\mu_{s,u}[k]$ using (38);
-5:         Update $\omega_{s,u}[k]$ using (39);
-6:         Update $\mathbf{W}_{s}$ by solving (IV-C3) via CVX;
-7:     until the reduction ratio of the objective value falls below a predefined threshold;
-8:end for
-9:Output: $\mathbf{W}_{s},\forall s$.
+With $\mu_{u}$ and $\omega_{u}$ fixed, the optimization problem to find the optimal $\mathbf{W}_{s}$ is formulated as
+$\mathop{\min}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\omega_{u}\Psi_{u}$
+${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{K},\;\forall s.$ (36)
+Here, $\Psi_{u}$ can be expressed as the sum of a positive semi-definite quadratic form and a linear term.
+We note that (IV-C3) is a convex QCQP with respect to $\mathbf{W}_{s}$, which can be solved using standard toolboxes such as CVX.
+However, this approach incurs a per-iteration complexity of $\mathcal{O}((SNU)^{3})$, since the beamformers of all satellites are optimized jointly.
+To better suit computation resource-constrained LEO scenarios, we instead develop a low-complexity method that strategically exploits BCD, strong duality, and matrix inversion properties.
+[^9] Despite the significant complexity reduction enabled by the proposed scheme, we note that the optimization is still carried out in a centralized manner (e.g., at a cluster-head satellite), which may incur scalability bottlenecks as the network size grows.
+Designing decentralized strategies that distribute the optimization across participating satellites to improve scalability is an important research direction but lies beyond the scope of this paper.
+We refer interested readers to our recent work [57] for a scalable distributed beamforming framework over networked LEO satellites.
+This reduces the dominant complexity of solving (IV-C3) from $\mathcal{O}((SNU)^{3})$ to $\mathcal{O}(SN^{3})$.
+Further details are provided in Appendix D.
+The overall procedure for solving (IV-B) is summarized in Algorithm 1, and its convergence follows directly from the BCD structure of the algorithm.
+$R_{u}=\Delta f\sum_{k=1}^{K}\log_{2}\!\left(1+\frac{\left|\sum\limits_{s=1}^{S}e^{-\jmath 2\pi k\Delta f\left(\tau_{s,u}-\tilde{\tau}_{s,u}\right)}\bar{\mathbf{h}}_{s,u}^{\mathsf{T}}\!\left[k\right]\mathbf{w}_{s,u}\right|^{2}}{\sum\limits_{m\neq u}^{U}\left|\sum\limits_{s=1}^{S}e^{-\jmath 2\pi k\Delta f\left(\tau_{s,u}-\tilde{\tau}_{s,m}\right)}\bar{\mathbf{h}}_{s,u}^{\mathsf{T}}\!\left[k\right]\mathbf{w}_{s,m}\right|^{2}+N_{0}\Delta f}\right),$ (38)
+IV-D Evaluation Model
+As stated above, after obtaining the cooperative beamformers by solving the optimization problem under the nominal signal model (with PACE-estimated channel parameters and ideal inter-satellite compensation), we evaluate the achievable communication performance using a more realistic evaluation model that employs the ground-truth channel parameters and explicitly accounts for practical compensation imperfections.
+This separation is intentional: it enables a direct assessment of the effectiveness of PACE by examining how well beamformers optimized with nominal (PACE-estimated) parameters perform when deployed in a realistic setting.
+Following the Doppler-compensation-free assumption adopted in[18, 24], we do not rely on explicit Doppler precompensation at the transmitter in the evaluation model for simplicity.
+Meanwhile, we account for imperfect delay compensation, which induces residual phase rotations across subcarriers.
+In addition, motivated by [58, 22, 23], we further incorporate the resulting asynchronous interference due to link-dependent delay mismatch across users in multi-LEO cooperative transmission.
+Specifically, after delay precompensation, the received signal of $u$-th UT on $k$-th subcarrier is modeled as [^10] Note that we assume the difference between $\tilde{\tau}_{s,m}$ and $\tau_{s,m}$ is smaller than the cyclic prefix (CP) length.
+$y_{u}\!\left[k\right]=\sum_{s=1}^{S}e^{-\jmath 2\pi k\Delta f\left(\tau_{s,u}-\tilde{\tau}_{s,u}\right)}\bar{\mathbf{h}}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,u}\,s_{u}\!\left[k\right]$ (37)
+$\quad+\sum_{m\neq u}^{U}\sum_{s=1}^{S}e^{-\jmath 2\pi k\Delta f\left(\tau_{s,u}-\tilde{\tau}_{s,m}\right)}\bar{\mathbf{h}}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,m}\,s_{m}\!\left[k\right]+z_{u}\!\left[k\right],$
+where $\bar{\mathbf{h}}_{s,u}[k]=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})\mathbf{a}\!\left(\bm{\theta}_{s,u}\right)$ denotes the (approximately) frequency-flat array response term, while the residual delay mismatch is explicitly captured by the subcarrier-dependent phase rotations in (37).
+Accordingly, the achievable rate of $u$-th UT under the evaluation model is given by (38) at the bottom of this page, which makes explicit that asynchronous interference persists due to link-dependent delay mismatch across users, even after applying delay precompensation.
+V Numerical Results
+V-A Simulation Parameters
+We model the Earth as a sphere with a radius of $6400\text{ km}$ and adopt a spherical coordinate system centered at the Earth’s core.
+The UTs are randomly distributed in a circular service area on the Earth’s surface with a radius of 200 $0\text{ km} [2$8].
+The participating LEO satellites are positioned within a corresponding circular area in space, centered along the extension of the line connecting the Earth’s core to the center of the surface service area, at an orbital altitude of $500\text{ km}$.
+Each satellite travels along an orbital path intersecting this line at a constant speed $\left|\mathbf{v}_{s}\right|=7.6\text{ km/s}$, with velocity vectors $\mathbf{v}_{s}$ tangent to their orbits and UPAs oriented toward Earth’s core.
+Due to imperfect time-domain and frequency-domain synchronization, the clock bias (respectively, CFO) between each satellite and the UT is randomly set to a value between $8\text{ ns}$ and $12\text{ ns}$ (respectively, between $\Delta f/120$ and $\Delta f/80$), which remains fixed throughout the simulation.
+To eliminate satellite handover involvement, we consider one large-timescale position-coherent frame containing $100$ small-timescale channel gain-coherent subframes, where the channel gain changes independently from one subframe to another.
+Throughout this frame, the satellite constellation topology remains constant for simplicity.
+The downlink positioning and uplink channel estimation consist of $L_{\text{p}}=10000$ and $L_{\text{c}}=1000$ pilot symbols, respectively.
+The channel gain from each LEO satellite to each UT, expressed by (3), is independently generated in each channel-gain coherent slot.
+The random phase $\psi_{s,u}$ is uniformly cooperative in $0,2\pi]$.
+The components that constitute the path loss $\beta_{s,u}$ are generated according to (4).
+The free-space path loss is given by $\beta^{\text{FS}}_{s,u}=20\log_{10}(\|\mathbf{p}_{u}-\mathbf{p}_{s}\|)+20\log_{10}(f_{c})-147.55\text{ [dB]}$[41].
+The atmospheric absorption $\beta^{\text{AB}}_{s,u}$, which depends on signal frequency and satellite elevation angle, is calculated according to International Telecommunication Union (ITU) recommendations [59] and configured using an off-the-shelf MATLAB script[36].
+In the considered LOS-dominant suburban/rural/remote scenarios, the clutter loss $\beta^{\text{CL}}_{s,u}$ and the shadow fading $\beta^{\text{SF}}_{s,u}$ originating from near-surface structures are typically weak[60].
+Therefore, we set $\beta^{\text{CL}}_{s,u}=0\text{ dB}$ and $\beta^{\text{SF}}_{s,u}=0\text{ dB}$ for convenience.
+Additionally, ionospheric scintillation can be ignored for frequencies above $6\text{ GHz}$[41].
+For $\beta^{\text{CL}}_{s,u}$ caused by tropospheric scintillation, which is generally difficult to model analytically, we adopt the example suggested in [36] as a reference value.
+The remaining simulation parameters are listed in Table I.
 TABLE I: Simulation Parameters
 Parameter Value
-[…]
-[…]
+Central carrier frequency $f_{c}$ $12.7\text{ GHz}$ (Ku band)
+Subcarrier spacing $\Delta f$ $120\text{ KHz}$
 Subcarrier number $K$ 1024
-[…]
-[…]
-[…]
-[…]
-Number of LEO satellites $S$ 8
-Number of UTs $U$ 4
-Antenna number $N=N_{\text{h}}\times N_{\text{v}}$ $8\times 8$
-Antenna radiation gain $G(\cdot)$ $\cos(\cdot)$
-V Numerical Results
-V-A Simulation Parameters
-Figure 4: Illustrative example of the considered scenario.
-[…]
-[…]
+Power budget at each LEO satellite $50\text{ dBm}$
+Power budget at each UT $40\text{ dBm}$
+PSD $N_{0}$ $173.855\text{ dBm/Hz}$
+Noise figure $F$ $10\text{ dB}$
+Number of LEO satellites $S$ 4
+Number of UTs $U$ 8
+Antenna number $N=N_{\text{h}}\times N_{\text{v}}$ $16\times 16$
+Antenna radiation gain $G(\cdot)$ $\sqrt{\frac{3}{4\pi}}\cos(\theta)$[39]
 V-B Compared Schemes
@@
 V-B3 Downlink Communication
-To evaluate the performance of downlink communication, we utilize the CSI obtained through the different channel estimation schemes introduced above to perform distributed beamforming, as summarized in Algorithm 1.
-The resulting beamformer is tested using ground-truth CSI, yielding the actual sum rate.
-Additionally, we consider two alternative benchmarks:
-  • PAB: This approach is identical to the one introduced for downlink positioning, which relies solely on position information.
-However, the beamformer at each LEO satellite is applied only to a subset of subcarriers.
-  • Single-Satellite Service ($\text{S}^{3}$): In this method, each UT is served by only a single LEO satellite.
-Specifically, the one providing the strongest signal strength, rather than being served by all satellites within a cluster.
+To evaluate downlink communication performance, we use the estimated CSI from the channel estimation schemes above to perform cooperative beamforming, as outlined in Algorithm 1 (since the satellites only have access to estimated CSI).
+The resulting beamformer is then tested with ground-truth CSI to obtain the actual sum rate.
+We also consider two benchmark schemes:
+  • PAB: This approach is the same as the one introduced for downlink positioning.
+  • No-Cooperation (NC): In this method, each UT is served only by the single LEO satellite with the strongest signal strength, thereby avoiding inter-satellite cooperation.
+The WMMSE algorithm is applied independently at each satellite to optimize its beamformers based solely on its own CSI for the associated UTs, while signals from other satellites are treated as interference.
 V-C Downlink Positioning
-$10^{-1}$$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$er-LEO satellite power budget [dBm]RMSE [m]PAB: LBPAB: MCRBPAB: BiasPAB: CRBVDB: LBVDB: MCRBVDB: BiasVDB: CRB
+[…]
 Figure 3: Positioning RMSE LB versus per-LEO satellite power budget.
 As shown in Fig.
-5, we compare the positioning error bounds and biases under PAB and VDB as functions of the per-satellite power budget.
-The results reveal that the LB of positioning error without model mismatch (represented by the positioning CRB under perfectly time- and frequency-synchronized satellites) decreases with increasing power.
-In contrast, the LB with mismatch initially decreases but eventually saturates at a value determined by the bias, which remains independent of the transmit power.
-This emphasizes the detrimental impact of imperfect time- and frequency-synchronization among LEO satellites on positioning performance [26, 24], potentially leading to overly optimistic performance estimates.
-Moreover, compared to VDB, PAB achieves a lower LB, with the performance gain being more pronounced at lower (and more practical) transmit power levels.
-This advantage arises because PAB, by utilizing position information, can allocate more power to the UTs than VDB, underscoring the importance of integrating prior position knowledge into the positioning system.
+3, we compare the positioning error bounds and biases under PAB and VDB with different per-satellite power budget.
+The results reveal that the LB of positioning error without model mismatch, represented by the positioning CRB under perfectly time- and frequency-synchronized satellites, decreases as the transmit power increases.
+In contrast, when synchronization mismatch is present, the LB initially decreases but eventually saturates due to the persistent positioning bias, which remains unaffected by the transmit power.
+This highlights the detrimental impact of imperfect time- and frequency-synchronization among LEO satellites on positioning performance [37, 35], potentially leading to overly optimistic performance evaluations.
+Furthermore, PAB consistently outperforms VDB by achieving a lower LB, with the performance gain being particularly significant (over an order of magnitude) at lower yet more practical transmit power levels.
+This advantage stems from PAB’s ability to leverage position information to allocate more power to the UTs than VDB, underscoring the importance of incorporating prior position knowledge into the design of positioning systems.
 […]
 Figure 4: Positioning RMSE LB versus maximum clock bias and CFO.
-As illustrated in Fig.
-6, we further analyze the impact of varying time and frequency biases among satellite-UT pairs on positioning performance, resulting from imperfect time and frequency synchronization among LEO satellites.
-Specifically, we present the LB of positioning error (under PAB) as a function of the maximum clock bias and CFO values.
-The clock bias and CFO for each satellite-UT pair are uniformly distributed between 00 and their respective maximum values.
-The results reveal that positioning performance deteriorates significantly as the maximum clock bias and CFO increase.
-A slight clock bias (on the order of milliseconds) or a minor CFO (on the order of $\Delta f/10$) can result in positioning errors of several kilometers, highlighting the critical importance of maintaining precise synchronization among LEO satellites in both time and frequency domains to achieve high-accuracy positioning.
-$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{-15}$$10^{-13}$$10^{-11}$$10^{-9}$osition error [m]RMSECRBLBMCRBBias
+In Fig.
+4, we analyze the impact of varying clock bias and CFO among satellite–UT pairs on the LB of positioning error (under PAB), where both impairments are assumed to be uniformly distributed between 0 and their respective maximum values.
+The results show that positioning performance degrades significantly as either factor increases.
+In particular, even a clock bias on the order of tens to hundreds of nanoseconds or a CFO of several kHz can lead to positioning errors exceeding $100\text{ m}$.
+These findings highlight the comparable severity of both impairments and underscore the critical importance of mitigating mismatches in both the time and frequency domains to ensure high-accuracy LEO positioning.
+$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{-6}$$10^{-4}$$^{-2}1001$$^{0}$Position error [m]Normalized RMSEPerfect-PACE: CRBPACE: LBPACE: MCRBPACE: Bias
 (a) Channel gain estimation
-$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{-13}$$10^{-11}$$10^{-9}$osition error [m]RMSE [s]
+$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$^{5}1010$$^{-10}$$10^{-8}$$10^{-6}$Position error [m]Normalized RMSE
 (b) TOA estimation
@@
 In Figs.
-7(a) and (b), we evaluate the LB for both channel gain and TOA estimation RMSE using our proposed PACE across varying positioning errors, comparing it with Perfect-PACE (without positioning error-induced model mismatch).
+5(a) and (b), we evaluate the normalized RMSE, obtained by dividing the RMSE by the ground-truth parameters, for both channel gain and TOA estimation using our proposed PACE, under varying positioning errors.
+The results are compared with those of Perfect-PACE, which assumes no model mismatch due to positioning errors.
 The results reveal that when positioning errors are small, PACE’s LB closely matches that of Perfect-PACE for both parameters.
 However, as positioning errors increase, PACE’s performance deteriorates due to growing model mismatch.
-For channel gain estimation, we observe that while the MCRB-induced error remains relatively insensitive to positioning errors, the bias component increases proportionally with positioning error, eventually becoming the primary contributor to overall channel gain error.
-The TOA estimation exhibits different behavior: its bias initially increases with positioning error before reaching a plateau, while the MCRB-induced error continues to grow, ultimately becoming the dominant factor in the total TOA estimation error.
-These findings underscore the importance of accurate downlink positioning in the first stage of PACE for achieving high-accuracy uplink channel estimation in the second stage.
+For channel gain estimation, we observe that while the MCRB-induced error remains relatively insensitive to positioning errors, the bias component increases proportionally with positioning error, eventually becoming the primary contributor to the overall channel gain error.
+The TOA estimation exhibits different behavior: its bias initially increases with positioning error before reaching a plateau, whereas the MCRB-induced error continues to grow, ultimately becoming the dominant factor in the total TOA estimation error.
+Nevertheless, the results indicate that channel estimation degrades noticeably only when the positioning error becomes impractically large (e.g., exceeding $1\text{ km}$).
+This observation highlights the robustness of PACE to positioning imperfections in LOS-dominant scenarios with relatively high satellite altitudes.
 In Figs.
-8(a) and (b), we compare the LB for the RMSE of both channel vector and TOA estimation using the proposed PACE against those obtained under UCE, versus UT’s power budget.
+6(a) and (b), we compare the LB for the RMSE of both channel vector and TOA estimation using the proposed PACE against those obtained under UCE, versus UT’s power budget.
 Note that PACE estimates both the channel gain $\alpha_{s,u}$ and the TOA $\tau_{s,u}$, whereas UCE estimates the channel vector $\bm{h}_{s,u}=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})\mathbf{a}(\bm{\theta}_{s,u})$ (excluding the TOA-induced phase) in addition to $\tau_{s,u}$.
 To enable a unified comparison, we transform the LB for channel gain estimation RMSE in PACE into the corresponding LB for channel vector estimation RMSE (using the form $\hat{\bm{h}}_{s,u}=\hat{\alpha}_{s,u}G(\hat{\theta}_{s,u}^{\mathrm{el}})\mathbf{a}(\hat{\bm{\theta}}_{s,u})$) through algebraic manipulation.
-As shown, the TOA estimation performance of PACE closely aligns with that of UCE, exhibiting only negligible degradation due to positioning error.
+As shown, the TOA estimation performance of PACE closely aligns with that of UCE, given that negligible positioning error is incurred in the positioning stage.
 This alignment occurs because TOA estimation is independent of whether the other unknown parameter is the channel gain or the channel vector.
@@
 Consequently, PACE achieves higher channel estimation accuracy under the same pilot length and power budget.
-[…]
-Figure 9: Sum rate versus per-LEO satellite antenna number.
-[…]
-Figure 10: Sum rate versus per-UT power budget.
 V-E Donwlink Communication
 In Fig.
-9, we compare the downlink communication sum rates under various approaches versus the number of antennas equipped on each LEO satellite.
-As observed, the sum rate under the proposed PACE significantly outperforms those under UCE and other baselines, reaching a level very close to the sum rate with perfect CSI.
-This results from PACE’s superior performance in reconstructing the channel accurately.
-Meanwhile, the sum rate under PAB is even lower than that under $\text{S}^{3}$, underscoring the importance of accurate CSI in delivering the full capability of multi-LEO satellite downlink communication.
-Furthermore, with an increasing number of antennas on each LEO satellite, the performance gain of the proposed scheme becomes more evident, highlighting its potential in facilitating high-speed LEO satellite communication with massive antenna arrays.
+7, we compare the downlink communication sum rates achieved by various approaches as a function of the number of antennas deployed on each LEO satellite.
+It is observed that Algorithm 1 combined with PACE significantly outperforms UCE and other baseline methods, achieving sum rates that closely approach those obtained with perfect CSI.
+This superior performance is attributed to PACE’s enhanced channel reconstruction accuracy and the benefits of multi-LEO satellite cooperation.
+In contrast, the sum rate under NC, whether using PACE or UCE, remains consistently higher than that achieved by PAB, even though PAB leverages inter-satellite cooperation.
+This highlights the critical role of accurate CSI in enhancing downlink communication efficiency, beyond what can be achieved using position information alone.
+Moreover, as the number of antennas per LEO satellite increases, the performance advantage of the proposed scheme, i.e., the combination of PACE and Algorithm 1, becomes increasingly pronounced, demonstrating its potential for enabling high-throughput LEO satellite communications with large-scale antenna arrays.
+4416166464256256055101015152020Antenna number at each LEO satelliteSum rate [Mbps]PACE + Algo. 1UCE + Algo. 1Perfect CSI + Algo. 1PABPACE + NCUCE + NC
+Figure 7: Sum rate versus per-LEO satellite antenna number.
+202030304040505060607070055101015152020Per-UT power budget [dBm]Sum rate [Mbps]PACE + Algo. 1UCE + Algo. 1Perfect CSI + Algo. 1PABPACE + NCUCE + NC
+Figure 8: Sum rate versus per-UT power budget.
 As shown in Fig.
-10, we compare the resultant sum rates under various schemes versus the per-UT power budget.
-It can be seen that the sum rates under PAB and $\text{S}^{3}$ are irrelevant to UTs’ transmit power, while $\text{S}^{3}$ remains superior to PAB throughout the entire range.
-However, with increasing transmit power from the UT side, the sum rates under PACE and UCE increase and gradually approach those under perfect CSI.
-This is because the channel estimation accuracy under both schemes increases with higher transmit power, which translates into a sum rate gain.
-Nevertheless, the sum rate under PACE is significantly always higher than that under UCE (especially notable when the transmit power at the UT is low).
-This highlights the importance of leveraging position information to obtain accurate CSI when the power budget at the UTs is limited, which is particularly relevant for equipments such as cell phones.
+8, we compare the sum rates under various schemes versus the per-UT power budget.
+It can be observed that the sum rate under PAB is independent of the UTs’ transmit power.
+In contrast, as the UTs’ transmit power increases, the sum rates achieved by Algorithm 1 and NC also increase.
+This is because higher transmit power improves channel estimation accuracy, which in turn reduces the mismatch between the channel parameters used to optimize beamforming and the ground-truth ones, thereby enhancing the resultant sum rate.
+Specifically, the sum rates achieved by Algorithm 1 significantly outperform those under NC and PAB, and gradually approach the performance achieved with perfect CSI.
+This demonstrates the superiority of multi-LEO satellite cooperative beamforming.
+Moreover, the use of PACE results in a substantially higher sum rate compared to UCE, especially when the UTs’ transmit power is low, due to its superior channel estimation accuracy.
+This highlights the importance of leveraging position information to obtain accurate CSI when the power budget at the UTs is limited, an aspect particularly relevant for power-constrained devices such as cell phones.
 VI Conclusion
@@
 Additionally, we theoretically demonstrate through MCRB analysis how position information improves uplink channel estimation even under inevitable positioning errors.
-To overcome the limited link budget offered by single-satellite-based downlink communication, we introduce a multi-satellite distributed beamforming strategy that relies on local CSI at each satellite, eliminating the need for centralized coordination.
+To overcome the limited link budget offered by single-satellite-based downlink communication, we introduce a multi-satellite cooperative beamforming optimization scheme to serve each UT simultaneously with multiple satellites within a cluster while reducing complexity.
 Numerical results validate our theoretical findings.
-As a direction for future work, developing practical algorithms and validating the framework through experimental testing would be valuable next steps.
-Appendix A Derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to $\mathbf{r}_{u}$
+Beyond the positioning-aided communication emphasized in this work, studying the reverse direction is equally important for future research.
+For example, inter-UT sidelink communication and ranging enable information exchange among proximal UTs and provide independent timing and distance references.
+Such mechanisms are promising for mitigating inter-satellite clock biases and thereby improving overall positioning accuracy.
+Appendix A MCRB Basics
+This subsection briefly recaps the fundamentals of MCRB.
+For a parameter $\bar{\bm{\theta}}\in\mathbb{R}^{M\times 1}$, the LB matrix of the mean squared error of a mismatched estimator is provided by[38]
+$\text{LBM}\big(\tilde{\bm{\theta}}\big)=\underbrace{\mathbf{A}^{-1}_{\tilde{\bm{\theta}}}\mathbf{B}_{\tilde{\bm{\theta}}}\mathbf{A}^{-1}_{\tilde{\bm{\theta}}}}_{\text{MCRB}\left(\tilde{\bm{\theta}}\right)}+\underbrace{\left(\bar{\bm{\theta}}-\tilde{\bm{\theta}}\right)\left(\bar{\bm{\theta}}-\tilde{\bm{\theta}}\right)^{\top}}_{\text{Bias}\left(\tilde{\bm{\theta}}\right)}$.
+(39)
+Here, $\tilde{\bm{\theta}}$ denotes the pseudo-true parameter obtained by
+$\tilde{\bm{\theta}}=\arg\underset{\bm{\theta}}{\min\;}\mathcal{D}\left\{\xi_{\text{T}}\left(\mathbf{y};\bar{\bm{\theta}}\right)\left|\right|\xi_{\text{M}}\left(\mathbf{y};\bm{\theta}\right)\right\},$ (40)
+[…]
+$\left[\mathbf{A}_{\tilde{\bm{\theta}}}\right]_{i,j}=$ $\mathbb{E}_{f_{T}}\left\{\frac{\partial^{2}\ln\xi_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}\partial\left[\bm{\theta}\right]_{j}}\bigg|_{\bm{\theta}=\tilde{\bm{\theta}}}\right\}$
+$=$ $2\Re\left\{\left(\frac{\partial^{2}\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}\partial\left[\bm{\theta}\right]_{j}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\bm{\epsilon}\left(\bm{\theta}\right)\right$.
+(41)
+$\;\;\;\;\;\;\;\;\;-\left.\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right)\right\}\bigg|_{\bm{\theta}=\tilde{\bm{\theta}}}$,
+and
+$\left[\mathbf{B}_{\tilde{\bm{\theta}}}\right]_{i,j}=$ $\mathbb{E}_{f_{T}}\left\{\frac{\partial\ln f_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\bigg|_{\bm{\theta}=\tilde{\bm{\theta}}}\frac{\partial\ln f_{M}\left(\mathbf{y};\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\bigg|_{\bm{\theta}=\tilde{\bm{\theta}}}\right\}$
+$=$ $4\Re\left\{\bm{\epsilon}\left(\bm{\theta}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right\}\Re\left\{\bm{\epsilon}\left(\bm{\theta}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right\}$
+$+$ $2\Re\left\{\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{i}}\right)^{\mathsf{H}}\mathbf{C}_{\text{M}}^{-1}\left(\frac{\partial\bm{\mu}\left(\bm{\theta}\right)}{\partial\left[\bm{\theta}\right]_{j}}\right)\right\}\bigg|_{\bm{\theta}=\tilde{\bm{\theta}}}$, (42)
+[…]
+Appendix B Derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to $\mathbf{r}_{u}$
 1) Non-Zero First-Order Derivative:
@@
 $\frac{\partial^{2}\upsilon_{s,u}}{\partial\mathbf{p}_{u}^{2}}$ $=-\frac{\mathbf{v}_{s}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}}$
-$-\frac{\mathbf{v}_{s}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}+\mathbf{v}_{s}^{\mathsf{T}}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\mathbf{I}_{3}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}},$
+$-\frac{\mathbf{v}_{s}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}+\mathbf{v}_{s}^{\mathsf{T}}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\mathbf{I}_{3}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}}$
 $+\frac{3\left(\mathbf{v}_{s}^{\mathsf{T}}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\right)\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{5}},$
 $\frac{\partial^{2}\tau_{s,u}}{\partial\mathbf{p}_{u}^{2}}$ $=\frac{\mathbf{I}_{3}}{c\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}-\frac{\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}}{c\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}}.$
-Appendix B Derivatives of $\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)$ with respect to $\bm{\eta}_{s,u}$
+Appendix C Derivatives of $\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)$ with respect to $\bm{\eta}_{s,u}$
 Note that the derivatives $\partial\tilde{\mathbf{g}}(\bm{\eta}_{s,u})/\partial[\bm{\eta}_{s,u}]_{i}$ and $\partial^{2}\tilde{\mathbf{g}}(\bm{\eta}_{s,u})/(\partial[\bm{\eta}_{s,u}]_{i}\partial[\bm{\eta}_{s,u}]_{j})$ can be constructed by concatenating $\partial\tilde{\mathbf{g}}_{s,u}[\ell,k]/\partial[\bm{\eta}_{s,u}]_{i}$ and $\partial^{2}\tilde{\mathbf{g}}_{s,u}[\ell,k]/(\partial[\bm{\eta}_{s,u}]_{i}\partial[\bm{\eta}_{s,u}]_{j})$, respectively, for $\ell=1,\dots,L_{\text{c}}$ and $k=1,\dots,K$.
@@
 $\frac{\partial^{2}\tilde{\mathbf{g}}_{s,u}\left[\ell,k\right]}{\partial\tau_{s,u}^{2}}$ $=-\vartheta_{s,u}(2\pi k\Delta f)^{2}e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\tilde{\mathbf{d}}_{s,u}\left[\ell,k\right].$
+Appendix D Low-Complexity Method for Solving (IV-C3)
+D-A Extracting Per-Satellite Subproblem
+To reduce complexity of solving (IV-C3), we instead adopt a BCD strategy, optimizing beamformers for each satellite sequentially while fixing the others.
+For the $s$-th satellite, the subproblem depends only on $\mathbf{w}_{s}=[\mathbf{w}_{s,1}^{\mathsf{T}},\ldots,\mathbf{w}_{s,U}^{\mathsf{T}}]^{\mathsf{T}}$ and can be reformulated as
+$\mathop{\min}\limits_{\mathbf{w}_{s}}\;\;\;$ $-2\sum_{u=1}^{U}\Re\left\{\omega_{u}\mu_{u}\mathbf{h}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}\right\}$
+$+\sum_{u=1}^{U}\sum_{m=1}^{U}\omega_{u}\left|\mu_{u}\right|^{2}\left|\mathbf{h}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}+\Omega_{s,u,m}\right|^{2}$
+${\rm{s.t.}}\;\;\;$ $\left\|\mathbf{w}_{s}\right\|^{2}\leq\frac{P_{s}}{K},$ (47)
+where $\Omega_{s,u,m}=\sum_{s^{{}^{\prime}}\neq s}^{S}\mathbf{h}_{s^{{}^{\prime}},u}^{\mathsf{T}}\mathbf{w}_{s^{{}^{\prime}},m}$.
+Let $\mathbf{E}_{u}=[\mathbf{0}_{N\times N},\ldots,\underbrace{\mathbf{I}_{N}}_{\text{The }u\text{-th matrix}},\ldots,\mathbf{0}_{N\times N}]\in\mathbb{R}^{N\times NU}$ be the selection matrix.
+Through algebraic manipulation, the above can be recast as a convex QCQP
+$\mathop{\min}\limits_{\mathbf{w}_{s}}\;\;\;$ $\mathbf{w}_{s}^{\mathsf{H}}\mathbf{Q}_{s}\mathbf{w}_{s}-2\Re\left\{\mathbf{b}_{s}^{\mathsf{H}}\mathbf{w}_{s}\right\}$
+${\rm{s.t.}}\;\;\;$ $\left\|\mathbf{w}_{s}\right\|^{2}\leq\frac{P_{s}}{K},$ (48)
+where
+$\mathbf{R}_{s}=\sum_{u=1}^{U}\omega_{u}\left|\mu_{u}\right|^{2}\mathbf{h}_{s,u}^{*}\mathbf{h}_{s,u}^{\mathsf{T}}\succeq\mathbf{0}_{N\times N},$ (49a)
+$\mathbf{Q}_{s}=\mathbf{I}_{U}\otimes\mathbf{R}_{s},$ (49b)
+$\mathbf{b}_{s}=\sum_{u=1}^{U}\omega_{u}\mu_{u}^{*}\mathbf{E}_{u}^{\mathsf{T}}\mathbf{h}_{s,u}^{*}-\sum_{u=1}^{U}\sum_{m=1}^{U}\omega_{u}\left|\mu_{u}\right|^{2}\Omega_{s,u,m}\mathbf{E}_{m}^{\mathsf{T}}\mathbf{h}_{s,u}^{*}.$
+D-B Exploiting Strong Duality
+It is straightforward to observe that the above problem is a convex QCQP involving only $\mathbf{w}_{s}$.
+Rather than relying on a convex solver such as CVX, which uses interior-point methods and incurs a complexity of $\mathcal{O}((NU)^{3})$, we exploit the problem’s strong duality to further reduce computational complexity.
+Specifically, the Lagrangian of the problem is
+$\mathcal{L}\left(\mathbf{w}_{s},\lambda\right)=\mathbf{w}_{s}^{\mathsf{H}}\mathbf{Q}_{s}\mathbf{w}_{s}-2\Re\left\{\mathbf{b}_{s}^{\mathsf{H}}\mathbf{w}_{s}\right\}-\lambda\left(\left\|\mathbf{w}_{s}\right\|^{2}-\frac{P_{s}}{K}\right),$
+where $\lambda\geq 0$ is the Lagrangian multiplier associated with the power constraint.
+Then, the optimal solutions of (D-A) can be derived through examining its Karush–Kuhn–Tucker (KKT) conditions as follows
+$\left(\mathbf{Q}_{s}+\lambda\mathbf{I}_{NU}\right)\mathbf{w}_{s}$ $=\mathbf{b}_{s},$ (50a)
+$\left\|\mathbf{w}_{s}\right\|$ $=\sqrt{\frac{P_{s}}{K}},$ (50b)
+$\lambda$ $\geq 0.$ (50c)
+It follows from (50a) that
+$\mathbf{w}_{s}=\left(\mathbf{Q}_{s}+\lambda\mathbf{I}_{NU}\right)^{-1}\mathbf{b}_{s}.$ (51)
+Therefore, we must choose $\lambda$ such that $\left\|(\mathbf{Q}_{s}+\lambda\mathbf{I}_{NU})^{-1}\mathbf{b}_{s}\right\|=\sqrt{P_{s}/K}$. Define
+$g\left(\lambda\right)=\left\|(\mathbf{Q}_{s}+\lambda\mathbf{I}_{NU})^{-1}\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}.$ (52)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 9: Run time and sum rate comparison between the QCQP-based and proposed low-complexity algorithms: (a) Run time versus LEO satellite number; (b) Sum rate versus LEO satellite number; (c) Run time versus antenna number; (d) Sum rate versus antenna number; (e) Run time versus UT number; (f) Sum rate versus UT number.
+D-C Reducing Matrix Inversion Complexity
+We note that (52) involves a $NU$-dimensional matrix inversion, which still incurs a computational complexity of $\mathcal{O}((NU)^{3})$ per iteration of the line search.
+To reduce this complexity, we exploit the structure $\mathbf{Q}_{s}=\mathbf{I}_{U}\otimes\mathbf{R}_{s}$.
+Let $\mathbf{b}_{s}=[\mathbf{b}_{s,1}^{\mathsf{T}},\ldots,\mathbf{b}_{s,U}^{\mathsf{T}}]^{\mathsf{T}}$, where $\mathbf{b}_{s,u}\in\mathbb{C}^{N}$.
+We then have
+$g\left(\lambda\right)$ $=\left\|\left(\mathbf{I}_{U}\otimes\mathbf{R}_{s}+\lambda\mathbf{I}_{NU}\right)^{-1}\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}$
+$=\left\|\left(\mathbf{I}_{U}\otimes\left(\mathbf{R}_{s}+\lambda\mathbf{I}_{N}\right)\right)^{-1}\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}$
+$=\left\|\left(\mathbf{I}_{U}\otimes\left(\mathbf{R}_{s}+\lambda\mathbf{I}_{N}\right)^{-1}\right)\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}$
+$=\left\|\left(\mathbf{I}_{U}\otimes\left(\mathbf{U}\left(\mathbf{\Lambda}+\lambda\mathbf{I}_{N}\right)^{-1}\mathbf{U}^{\mathsf{H}}\right)\right)\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}$
+$=\sum_{u=1}^{U}\left\|\mathbf{U}\left(\mathbf{\Lambda}+\lambda\mathbf{I}_{N}\right)^{-1}\mathbf{U}^{\mathsf{H}}\mathbf{b}_{s,u}\right\|^{2}-\frac{P_{s}}{K}$
+$=\sum_{u=1}^{U}\sum_{n=1}^{N}\frac{\left|\varpi_{s,u,n}\right|^{2}}{\left(\lambda_{n}+\lambda\right)^{2}}-\frac{P_{s}}{K},$ (53)
+where $\bm{\varpi}_{s,u}=[\varpi_{s,u,1},\ldots,\varpi_{s,u,N}]^{\mathsf{T}}=\mathbf{U}^{\mathsf{H}}\mathbf{b}_{s,u}$.
+Here, $\mathbf{R}_{s}$ is decomposed as $\mathbf{R}_{s}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\mathsf{H}}$ via eigenvalue decomposition, where $\mathbf{U}\in\mathbb{C}^{N\times N}$ is a unitary matrix and $\mathbf{\Lambda}=\text{diag}[\lambda_{1},\ldots,\lambda_{N}]$ is the diagonal eigenvalue matrix.
+Then, (51) is reformulated as
+$\mathbf{w}_{s}=\left(\mathbf{I}_{U}\otimes\left(\mathbf{U}\left(\mathbf{\Lambda}+\lambda\mathbf{I}_{N}\right)^{-1}\mathbf{U}^{\mathsf{H}}\right)\right)\mathbf{b}_{s},$ (54)
+which eliminates the need for an $NU$-dimensional matrix inversion and thereby reduces the computational complexity of updating beamformers.
+D-D Analyzing Complexity
+The above reformulation shows that eigenvalue decomposition of the matrix $\mathbf{R}_{s}$ needs to be performed only once with complexity $\mathcal{O}(N^{3})$, rather than repeatedly inverting an $NU$-dimensional matrix for each $\lambda$.
+The Golden-section search method, which has a linear convergence rate, finds an $\epsilon$-solution within $\mathcal{O}(\log(1/\epsilon))$ iterations.
+Since each iteration only requires evaluating a one-dimensional function, the associated complexity can generally be neglected.
+Consequently, the eigenvalue decomposition step dominates the overall complexity of solving (D-A).
+As (D-A) is solved sequentially for $s=1,\ldots,S$, the total complexity of solving (IV-C3) is reduced to $\mathcal{O}(SN^{3})$, compared with the original $\mathcal{O}((SNU)^{3})$.
+To further demonstrate the superiority of the proposed algorithm, Fig.
+9 compares the run time and sum rate of the QCQP-based and proposed low-complexity algorithms under varying numbers of LEO satellites, antennas per LEO, and UTs.
+The results show that the proposed algorithm, which decomposes the beamformer update subproblem via BCD and exploits duality and line search, achieves orders-of-magnitude complexity reduction while maintaining identical sum-rate performance compared with the original QCQP-based update.
+This is notable since the QCQP formulation yields the optimal subproblem solution, whereas the proposed BCD-based structure converges only to a stationary point.
+These results validate the efficiency of the proposed low-complexity design.